### v3

# Asymptotically Unbiased Synthetic Control Methods by Distribution Matching

Abstract
Synthetic Control Methods (SCMs) have become an essential tool for comparative case studies. The fundamental idea of SCMs is to estimate the counterfactual outcomes of a treated unit using a weighted sum of the observed outcomes of untreated units. The accuracy of the synthetic control (SC) is critical for evaluating the treatment effect of a policy intervention; therefore, the estimation of SC weights has been the focus of extensive research. In this study, we first point out that existing SCMs suffer from an endogeneity problem, the correlation between the outcomes of untreated units and the error term of the synthetic control, which yields a bias in the treatment effect estimator. We then propose a novel SCM based on density matching, assuming that the density of outcomes of the treated unit can be approximated by a weighted average of the joint density of untreated units (i.e., a mixture model). Based on this assumption, we estimate SC weights by matching the moments of treated outcomes with the weighted sum of moments of untreated outcomes. Our proposed method has three advantages over existing methods: first, our estimator is asymptotically unbiased under the assumption of the mixture model; second, due to the asymptotic unbiasedness, we can reduce the mean squared error in counterfactual predictions; third, our method generates full densities of the treatment effect, not merely expected values, which broadens the applicability of SCMs. We provide experimental results to demonstrate the effectiveness of our proposed method. [^1] This study was presented at the Workshop on Counterfactuals in Minds and Machines at the International Conference on Machine Learning in July 2023 and at the International Conference on Econometrics and Statistics in August 2023.
1 Introduction
Synthetic control methods (SCMs, Abadie and Gardeazabal, 2003; Abadie et al., 2010) have gained attention as a crucial tool for program evaluation in comparative case studies across fields such as economics, statistics, and machine learning. According to Athey and Imbens (2017), SCMs are “arguably the most important innovation in the policy evaluation literature in the last 15 years” due to their “simplicity of the idea, and the obvious improvement over standard methods.”
SCMs posit a scenario in which multiple units exist, with one unit subjected to a policy intervention (treated unit) at a certain time period, while the rest remain untreated. The focus is on estimating the treatment effect of the treated unit, i.e., the difference between its observed factual outcome and unobserved counterfactual outcome. Untreated units are used to account for unobserved trends in the outcome over time that are unrelated to the effect of the policy intervention. The key insight is that an optimally weighted average of the untreated units, referred to as the synthetic control (SC) unit, provides a more appropriate estimate of the counterfactual than estimating the counterfactual using only the treated unit. See Figure 1 for an illustration. SCMs have been applied in a variety of settings, such as studying the effect of terrorism on GDP (Abadie and Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham and Shah, 2017).
Figure 1: Illustration of SCMs
For instance, Abadie et al. (2010) used SCM to study the treatment effect of a comprehensive tobacco control program implemented in California in 1988. Smoking rates in California decreased following the program’s implementation, but it was unclear if this decrease was a result of the program or a long-term trend. To investigate the treatment effect, annual per-capita cigarette sales data across several states that did not implement a similar tobacco control program were used as untreated units. They then applied SCM to estimate California’s counterfactual outcome: the smoking rate in California in 1989 had the tobacco control program not been implemented. For further details, also see Abadie (2021).
A significant challenge in SCMs is the typical assumption that the expected outcome of the treated unit can be represented by the weighted sum of untreated units. However, Ferman and Pinto (2021) points out that there exists a non-negligible asymptotic bias in the obtained estimator. According to Ferman and Pinto (2021), this asymptotic bias is caused by a kind of endogeneity in the linear regression model deduced from the assumption, due to a correlation between the outcomes of untreated units and the error term of the linear model. This problem is akin to endogeneity caused by measurement error (Greene, 2003). We refer to this endogeneity as implicit endogeneity.
To address the aforementioned issue, we propose to assume a mixture model instead of a linear relationship among expected outcomes. Specifically, we assume that the density of the outcome of the treated unit can be approximated by a linear combination of those of the untreated units. This mixture model assumption also implies that each moment of the outcome of the treated unit can be represented by a linear combination of those of the untreated units. Therefore, we can estimate the SC weights by matching all moments between the outcome of the treated unit and those of the untreated units. We refer to our method as the Density Matching SCM (DMSCM). Unlike existing methods, our proposed estimator is asymptotically unbiased. We validate the efficacy of our proposed method through simulation studies and empirical analysis.
Our proposed method corresponds to a simplification of the distributional SCMs by Gunsilius (2023). Although Gunsilius (2023) does not explicitly point out the implicit endogeneity problem, it observes that their result does not coincide with the result under the standard SCMs. This is because under the mixture model assumption, while Gunsilius (2023) returns a consistent estimator for the treatment effect, existing methods do not. We connect the distributional SCMs to the asymptotic bias problem and propose a simplified estimation method for SCMs by distribution matching.
In summary, our contributions are as follows:
  • Reconsidering the assumptions necessary for treatment effect estimation in SCMs.
  • Proposing a density-matching method for asymptotically unbiased SCMs.
  • Simplifying the density-matching method.
1.1 Related Work
Following the pioneering work by Abadie and Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015). These SCMs, including their variants such as those by Doudchenko and Imbens (2016) and Ben-Michael et al. (2019), are based on minimizing the pre-treatment sum of squared errors between the outcome of the treated unit and the weighted sum of those of the untreated units, with some regularization and constraints regarding the weights.
Gunsilius (2023) explores distribution SCMs (DiSCo), which estimate SC weights to match the distribution of the outcomes of the treated unit and the weighted sum of the distributions of the untreated units. [^2] The method of Gunsilius (2023) is called DiSCo by Dijcke (2024). His method is based on minimizing the Wasserstein distance, which can be simplified as matching the quantiles in the setting of SCMs. Their assumption about identification can be interpreted as a variant of mixture models.
From a theoretical viewpoint, Abadie et al. (2010) demonstrate that, if there are weights that provide a perfect pre-treatment fit, then the bias of the SC estimator is bounded by a function that converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases, even when the number of untreated units $J$ is fixed. However, Ferman and Pinto (2021) finds that if the pre-treatment fit is imperfect, then the conventional SCMs do not yield a consistent estimator of the treatment effect due to the bias caused by the correlation between the outcomes of the untreated units and the error term in the posited linear models. We review this problem in Section 4 and name it as the implicit endogeneity.
To address this bias issue, existing studies have proposed several approaches. Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where the number of untreated units increases ($J\to\infty$) as well as the length of the pre-intervention periods $T_{0}$ also increases.
The use of mixture models has a close relationship with linear factor models, assumed in existing studies (Abadie, 2021). We justify the use of the former by employing arguments from Shi et al. (2022) and Nazaret et al. (2023). Our proposed method can be considered a new interpretation of existing SCMs. By reinterpreting existing SCMs through density matching, in addition to addressing the implicit endogeneity problem, we address two problems in the SCM literature. First, our proposed method improves the pre-treatment fit by debiasing the estimator. Second, our method generates full counterfactual densities of the treated unit, rather than merely the expected value. As a result, we can conduct statistical inference on the distribution, or any function of the distribution, such as quantiles (Gunsilius, 2023), Lorenz curves (Gastwirth, 1971), and expected social welfare (Abadie, 2002).
An excellent illustration of the relevance of estimating distribution functions is Ropponen (2011), who applies the changes-in-changes estimator by Athey and Imbens (2006) to estimate the effect of minimum wage changes on employment levels in order to reconcile the classical conflicting results about the minimum wage debate in Card and Krueger (2000) and Neumark and Wascher (2000). Spiess et al. (2023) investigates high-dimensional linear regression models in SCMs.
The exploration of identification assumptions has been attempted in the existing literature. Ferman and Pinto (2021) conducts a comprehensive investigation to discern the conditions under which SCMs can accurately predict counterfactual outcomes.
This study also answers an important open question raised by Gunsilius (2023). Gunsilius (2023) discusses that an optimal weight obtained by the density matching differs from that obtained by the classical least squares approach. Gunsilius (2023) conjectures that the classical least squares approach estimates a weight to match the first moment, while the density matching approach estimates a weight to match all moments. However, we point out that the classical least squares approach estimates a weight not only to match the first moment but also tends to incur bias due to the implicit endogeneity.
Concerning the assumptions about SCMs, Wan et al. (2018) points out that the SCMs essentially assume mixture models, which distinguish SCMs from traditional panel data analysis.
On another note, Shi et al. (2022) introduces fine-grained models to investigate hidden assumptions inherent to SCMs, which also explores a connection between the linear combination of expected outcomes and the linear factor models. Nazaret et al. (2023) extends these findings into the realm of model misspecification. Their formulation is also based on the mixture models.
Independently of us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias pointed out by Ferman and Pinto (2021). [^3] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and we and he conducted our studies completely independently. While our and his formulations about the endogeneity are similar, the approaches are significantly different. We address the problem from the viewpoint of density matching, which is an approach similar to Gunsilius (2023). In contrast, Fry (2024) uses an instrumental variable method by employing outcomes of untreated units that are not used in SCMs as instrumental variables.
Table 1: Comparison of SCMs. See the following sections for the definitions of variables and methods.
Method Assumption Asymptotic unbiasedness Remark
Ours DMSCM $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\ \ \ \forall y\in\mathbb{R}$ Unbiased
Abadie and Gardeazabal (2003) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Biased
Ferman (2021) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Unbiased $J\to\infty$
Gunsilius (2023) DiSCo $F^{-1}_{Y^{N}_{0,t}}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y^{N}_{j,t}}(q)\ \ \ \forall q\in(0,1)$ Unbiased
Fry (2024) IV method (GMM) $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Unbiased Use IVs
1.2 Organizations
In Section 2, we formulate the problem. In Section 3, we review an idea of SCMs. In Section 4, we point out the problem of implicit endogeneity. Then, in Section 5, we propose our method under the assumption of mixture models. In Section 6, we investiate the asymptotic property of our proposed method. In Section 7, we present a statistical inference method for our proposed method. In Section 8, we discuss related topics and existing studies. We numerically investigate the performance of the proposed estimators through extensive simulations in Section 9.
2 Problem Setting
Suppose that there are $J+1$ units, indexed by $j\in\mathcal{J}\coloneqq\{0,1,2,\dots,J\}$, and a time series indexed by $t\in\mathcal{T}=\{1,2,\dots,T\}$ where $3\leq T\in\mathbb{N}$. Without loss of generality, we assume that the first unit ($j=0$) is the treated unit, i.e., the unit affected by the policy intervention of interest. A set of potential comparisons, $j\in\mathcal{J}^{U}\coloneqq\mathcal{J}\backslash\{0\}=\{1,2,\dots,J\}$, consists of untreated units not affected by the intervention. We also assume that our data span $T$ periods and that the intervention occurs at $t=T_{0}$, such that $2\leq T_{0}<T$. Observations at period $t\in\mathcal{T}_{0}\coloneqq\{1,2,\dots,T_{0}\}$ are from before the intervention, and observations at period $t\in\mathcal{T}_{1}\coloneqq\mathcal{T}\backslash\mathcal{T}_{0}$ are from after the intervention. Let $T_{1}\coloneqq|\mathcal{T}_{1}|=T-T_{0}$.
2.1 Potential Outcomes
First, following the Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974), we define the potential outcomes of the policy intervention. For each unit $j\in\mathcal{J}$ and period $t\in\mathcal{T}$, there are two outcomes of interest, $\left(Y^{I}_{j,t},Y^{N}_{j,t}\right)\in\mathbb{R}^{2}$, corresponding to scenarios with and without the intervention. For each $j\in\mathcal{J}$ and $t\in\mathcal{T}$, let $F^{I}_{j,t}(y)$ and $F^{N}_{j,t}(y)$ be the cumulative distribution functions (CDFs) for $Y^{I}_{j,t}$ and $Y^{N}_{j,t}$, respectively.
2.2 Observations
We then define our observations. For each unit $j\in\mathcal{J}$, we can only observe one of the outcomes corresponding to the actual intervention; that is, we observe $Y_{j,t}\in\mathbb{R}$ such that
$Y_{0,t}=\begin{cases}Y^{N}_{0,t}&\forall t\in\mathcal{T}_{0}\\ Y^{I}_{0,t}&\forall t\in\mathcal{T}_{1}\end{cases},\qquad\mathrm{and}\qquad Y_{j,t}=Y^{N}_{j,t}\qquad\forall j\in\mathcal{J}^{U},\forall t\in\mathcal{T}.$
We aim to estimate the effect of the policy intervention by predicting $Y^{I}_{0,t}$ from $(Y^{N}_{j,t})_{j\in\mathcal{J}^{U}}$ and by comparing this prediction to the observed $Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$. We will define the treatment effect more formally in the following section.
2.3 Treatment Effect
In SCMs, there are several ways to define the treatment effect $\tau_{0,t}\in\mathbb{R}$ for each $t\in\mathcal{T}_{1}$ of the intervention for the treated unit. First, for each $t\in\mathcal{T}_{1}$, we define the treatment effect on the treated unit $j=0$ as
$\tau_{0,t}\coloneqq\mathbb{E}\big{[}Y^{I}_{0,t}\big{]}-\mathbb{E}\big{[}Y^{N}_{0,t}\big{]}.$
Another definition is based on the following linear factor model:
Definition 2.1.
We say that $(Y^{I}_{j,t},Y^{N}_{j,t})$ follows the linear factor model if:
$Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t},\quad\mathbb{E}_{j,t}[\epsilon_{j,t}]=0,\quad\forall t\in\mathcal{T},\quad\forall j\in\mathcal{J},$ (1)
$Y^{I}_{j,t}=\tau_{j,t}+Y^{N}_{j,t}.$
where $\delta_{t}$ is an unobserved common factor with constant factor loadings across units, $c_{j}$ is an unknown time-invariant fixed effect, $\lambda_{t}$ is a $(1\times F)$ vector of unobserved common factors, $\mu_{j}$ is a $(F\times 1)$ vector of unknown factor loadings, and the error terms $\epsilon_{j,t}$ are unobserved idiosyncratic shocks. Here, while $c_{j}$, $\mu_{j}$, and $\tau_{0,t}$ are non-random variables, $\delta_{t}$, $\lambda_{t}$, and $\epsilon_{j,t}$ are random variables.
Under the linear factor model, the treatment effect is given by $\tau_{0,t}=Y^{I}_{0,t}-Y^{N}_{0,t}$, a non-random variable by definition of the factor model. Therefore, the expression $\tau_{0,t}\coloneqq\mathbb{E}\big{[}Y^{I}_{0,t}\big{]}-\mathbb{E}\big{[}Y^{N}_{0,t}\big{]}$ also holds under the factor model.
Comparative case studies aim to evaluate $\tau_{0,t}$ by predicting $Y^{N}_{0,t}$ for $t>T_{0}$, which is the counterfactual outcome that would have been observed for the treated unit in the absence of the intervention. By definition, $Y^{N}_{0,t}$ is a counterfactual outcome because of the intervention after $t=T_{0}$. We usually use one unit or a small number of units that have similar characteristics to the treated unit at the time of the intervention but are not exposed to the intervention. However, when the data consist of a few units, such as regions or countries, it is often difficult to find a single untreated unit that provides a suitable comparison for the unit treated by the policy intervention of interest.
Distributional Treatment Effects. In program evaluation, we are often interested not only in the scalar treatment effect but also in the distributions of outcomes (Maier, 2011) or their functionals, such as the quantile treatment effects (QTEs) and expected social welfare (Abadie, 2002). In this study, we consider estimating $p^{I}_{0,t}(y)$ and $p^{N}_{0,t}(y)$ for $t\in\mathcal{T}_{1}$, where outcomes corresponding to $p^{I}_{0,t}(y)$ are unobservable. The treatment effect is then captured by $p^{I}_{0,t}(y)$ and $p^{N}_{0,t}(y)$, or functionals of these distributions, such as QTEs and expected social welfare. Depending on the application, we may adjust the target density, for instance, $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$. We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus and Oprescu, 2022; Chikahara et al., 2022). In the context of SCMs, Gunsilius (2023) and Chen (2020) have estimated the treatment effect.
3 Recap of SCMs
This section reviews the original SCM proposed by Abadie and Gardeazabal (2003). To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ by using the weighted sum of $\{Y^{N}_{j,t}\}^{J}_{j=1}$ for $t\in\mathcal{T}_{1}$; that is, SCMs predict $Y^{N}_{0,t}$ as
$\widehat{Y}^{N}_{0,t}\coloneqq w_{0}+\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t},$
where $w_{0}\in\mathbb{R}$ and $w_{j}\in\mathbb{R}$ $(j\in\mathcal{J}^{U})$ are the weights. We refer to these weights as SC weights.
Several methods exist for estimating the SC weights. In the SCM proposed by Abadie and Gardeazabal (2003), SC weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1t},\dots,Y^{N}_{J,t}$. Let us define
$\Delta^{J}\coloneqq\{\bm{w}=(w_{1}\ w_{2}\ \dots\ w_{J})^{\top}\in[0,1]^{J}\mid\sum_{j\in\mathcal{J}^{U}}w_{j}=1\}.$
Then, Abadie and Gardeazabal (2003) estimates the SC weights as
$\bm{w}^{\mathrm{Abadie}}=\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t}\right)^{2},$
The counterfactual outcome is then predicted as
$\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}=\widehat{w}^{\mathrm{Abadie}}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{Abadie}}_{j}Y^{N}_{j,t}.$
Using the prediction $\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$, we estimate the treatment effect as $\widehat{\tau}^{\mathrm{Abadie}}_{t}=Y^{I}_{0,t}-\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$.
The literature on inference for SCMs has grown significantly, extending beyond the original proposals of Abadie and Gardeazabal (2003) and Abadie et al. (2010). Notable contributions include works by Li (2020), Shaikh and Toulis (2021), Cattaneo et al. (2021), and Chernozhukov et al. (2019). Doudchenko and Imbens (2016) proposes estimating the SC weight as
$\widetilde{\bm{w}}^{\mathrm{DI}}=\operatorname*{arg\,min}_{\widetilde{\bm{w}}=(w_{0}\ w_{1}\ \cdots\ w_{J})^{\top}\in\mathbb{R}^{J+1}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-w_{0}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y_{j,t}\right)^{2}+\lambda_{1}\|\widetilde{\bm{w}}\|_{1}+\lambda_{2}\|\widetilde{\bm{w}}\|_{2},$
where $\lambda_{1}>0$ and $\lambda_{2}>0$ are regularization parameters. The parameter space for $\widetilde{\bm{w}}$ can be restricted in various ways, such as setting $w_{0}=0$ and $\bm{w}=(w_{1}\ \cdots\ w_{J})^{\top}\in\Delta^{J}$.
4 Identification and Endogeneity
This section addresses the identification problem of treatment effects from the perspective of endogeneity.
4.1 Linear Models in Expected Outcomes
We reconsider the underlying assumptions in SCMs. Existing studies often posit some linear relationship between $Y^{N}_{0,t}$ and $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. One such relationship assumes the existence of $\bm{w}^{*}\in\Delta^{J}$ (Abadie et al., 2010):
$Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}.$
This assumption, known as the perfect pre-treatment fit (Ferman and Pinto, 2021), implies that the stochastic process of $Y^{N}_{0,1},\dots,Y^{N}_{0,T}$ can be completely replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$. In linear factor models, $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, and this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is considered unrealistic (note that this is a specific case and not a necessary condition for $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$).
A more natural assumption is that each term in the linear factor models of $Y^{N}_{0,t}$ can be replicated by the weighted sum of corresponding terms in the models of $Y^{N}_{0,1},\dots,Y^{N}_{0,T}$. A weaker assumption still is a linear relationship between $\mathbb{E}_{0,t}\left[Y^{N}_{0,t}\right]$ and $(\mathbb{E}_{j,t}\left[Y^{N}_{j,t}\right])_{j\in\mathcal{J}^{U}}$, which includes the replication of each term in linear factor models as a specific case.
Assumption 4.1.
There exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$, it holds that
$\mathbb{E}_{0,t}\left[Y^{N}_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y^{N}_{j,t}\right],$ (2)
where $\Delta^{J}\coloneqq\{(w_{1},\dots,w_{J})\in\mathbb{R}^{J}\mid w_{j}\geq 0\ \forall j\ \mathrm{and}\ \sum_{j\in\mathcal{J}^{U}}w_{j}=1\}$.
Compared to Assumption 4.1, assuming $Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}Y_{j,t}$ is a stronger and less convincing assumption than (2) since $Y_{j,t}$ is a random variable. Therefore, in this study, we discuss SCMs based on (2). Ferman and Pinto (2021) explicitly considers Assumption 4.1.
4.2 Asymptotic Bias in Linear Factor Models
We consider the linear factor model as defined in Definition 2.1, along with Assumption 4.1. However, under the linear factor model, we adopt the following more specific assumption:
Assumption 4.2.
Under the linear factor model from Definition 2.1, there exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$, the following conditions hold:
$c_{0}=\sum_{j\in[J]}w^{*}_{j}c_{j}\quad\text{and}\quad\mu_{0}=\sum_{j\in[J]}w^{*}_{j}\mu_{j}\ \ \forall t\in\mathcal{T},\ \forall j\in\mathcal{J}.$
Assumption 4.2 is a specific case of Assumption 4.1; that is, if Assumption 4.1 holds, then Assumption 4.2 also holds under the linear factor model defined in Definition 2.1.
Under the linear factor model with Assumption 4.2, Ferman and Pinto (2021) finds that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{t}$ is biased and not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero. We further make the following assumption for the linear factor models.
Assumption 4.3 (Common and idiosyncratic shocks).
For each $j\in\mathcal{J}$, the linear factor models 2.1 satisfy the following conditions as $T_{0}\to\infty$:
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\lambda_{t}\xrightarrow{\mathrm{p}}0$.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\xrightarrow{\mathrm{p}}0$.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\lambda^{\top}_{t}\lambda_{t}\xrightarrow{\mathrm{p}}\Omega_{0}$, where $\Omega_{0}$ is positive semidefinite.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\epsilon^{\top}_{j,t}\xrightarrow{\mathrm{p}}\sigma^{2}_{j,\epsilon}I_{J+1}$, where $\sigma^{2}_{j,\epsilon}\geq 0$ is constant.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\lambda_{t}\xrightarrow{\mathrm{p}}0$.
Ferman and Pinto (2021) then demonstrates that the SCM proposed by Abadie and Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$ under these conditions.
Proposition 4.4 (Proposition 1 in Ferman and Pinto (2021)).
Under Assumptions 4.34.2, it holds that
$\bm{w}^{\mathrm{Abadie}}\xrightarrow{\mathrm{p}}\overline{\bm{w}}^{\mathrm{Abadie}}\coloneqq\left(\overline{w}^{\mathrm{Abadie}}_{j}\right)_{j\in\mathcal{J}^{U}}$
as $T_{0}\to\infty$, where $(c_{0},\mu_{0})\neq(\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j},\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j})$, unless $\sigma^{2}_{\epsilon}=0$ or $\widetilde{\Phi}\cap\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\{\bm{w}^{\top}\bm{w}\}\neq\emptyset$. Moreover, for $t\in\mathcal{T}_{1}$, it holds that
$\widehat{\tau}^{\mathrm{Abadie}}_{0,t}$ $\xrightarrow{\mathrm{p}}\tau_{0,t}+\lambda_{t}\left(\mu_{0}-\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right)+\left(c_{0}-\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right)+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}\epsilon_{j,t}\overline{w}^{\mathrm{Abadie}}_{j}\right)$
as $T_{0}\to\infty$.
To mitigate this issue, Ferman and Pinto (2021) proposes the demeaned SC estimator
$\widehat{\bm{w}}^{\mathrm{FP}}=\operatorname*{arg\,min}_{\widehat{\bm{w}}^{\mathrm{FP}}\in\Delta^{J}}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left\{\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\right\}^{2},$
where $\overline{Y}^{N}_{j}=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}Y_{j,t}$. They estimate $\tau_{0,t}$ as $\widehat{\tau}^{\mathrm{FP}}_{0,t}=Y^{I}_{0,t}-\overline{Y}^{N}_{0}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{FP}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)$. They show the following result.
4.3 Implicit Endogeneity
In this section, we interpret the asymptotic bias identified by Ferman and Pinto (2021) from the perspective of endogeneity.
To clarify this issue, we consider a simpler case using a least-squares estimator rather than $\bm{w}^{\mathrm{Abadie}}$:
$\widehat{\bm{w}}^{\mathrm{LS}}=\operatorname*{arg\,min}_{\bm{w}\in\mathbb{R}^{J}}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y_{j,t}\right)^{2}.$
We demonstrate that $\widehat{\bm{w}}^{\mathrm{LS}}$ is asymptotically biased due to endogeneity. Define $\epsilon_{j,t}=Y_{j,t}-\mathbb{E}_{j,t}\left[Y_{j,t}\right]$. Assuming that (2) holds, the following linear regression model is applicable:
$Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}+\nu_{t},\quad\text{where}\quad\nu_{t}=-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}+\epsilon_{0,t}.$
Here, $Y_{0,t}=\mathbb{E}_{0,t}[Y_{0,t}]+\epsilon_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]+\epsilon_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}+\epsilon_{0,t}$. Note that $\mathbb{E}\left[\nu_{t}Y_{j,t}\right]\neq 0$, implying endogeneity due to a correlation between the regressor and the error term. Endogeneity is known to introduce bias in the least-squares estimator. Define $Q^{*}$ as a $(J\times J)$-matrix whose $(i,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\mathbb{E}_{i,t}[Y_{i,t}]\mathbb{E}_{j,t}[Y_{j,t}]$. Let $\Sigma$ be a $J\times J$ diagonal matrix whose $(j,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\sigma^{2}_{j,t}$.
Theorem 4.5.
As $T_{0}\to\infty$, $\widehat{\bm{w}}^{\mathrm{LS}}\xrightarrow{\mathrm{p}}\overline{\bm{w}}^{\mathrm{LS}}$, where
$\overline{\bm{w}}^{\mathrm{LS}}=\bm{w}^{*}-\left(Q^{*}+\Sigma\right)^{-1}\Sigma\bm{w}^{*}.$
This theorem is derived from Section 5.6 of Greene (2003). The result implies that under the assumption of (2), there is an asymptotic bias in $\widehat{\bm{w}}^{\mathrm{LS}}$; that is, least-squares-type estimators do not converge to the true value $\bm{w}^{*}$.
This section highlights the identification of treatment effects and the problem of implicit endogeneity. Though the estimator $\bm{w}^{\mathrm{Abadie}}$ is widely used in existing studies, we find that it suffers from implicit endogeneity. This issue has been explicitly or implicitly discussed in previous studies (Chernozhukov et al., 2019; Ferman and Pinto, 2021; Ferman, 2021). Independently of our work, Fry (2024) explicitly addresses the bias as an endogeneity problem and applies the instrumental variable method to resolve the issue.
5 SCM by Density Matching
As discussed in Section 4, Assumption 4.1 is considered weak or incompatible for identification of $\bm{w}^{*}$ using least squares. Therefore, we consider adopting another assumption for identification. Specifically, we propose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$ and introduce a SCM based on the density-matching approach.
5.1 Identification and Estimation Strategy
Section 4 and the results in Ferman and Pinto (2021) imply that combining the assumption $\mathbb{E}_{0,t}\left[Y_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]$ with least squares estimators does not yield consistent estimators for $\bm{w}^{*}$. Hence, we consider another more realistic and theoretically favorable assumption.
We begin with the premise that the distribution of the target unit’s outcome, $Y^{N}_{0,t}$, is a linear combination of those of the control units’ outcomes, $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. For simplicity, in this subsection, we assume that $Y^{N}_{j,t}$ is a continuous random variable with a probability density function (PDF). We then consider a scenario where the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$:
$p^{N}_{0,t}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y)\qquad\forall t\in\mathcal{T}.$ (3)
We discuss how to estimate SC weights that satisfy (3). One approach is to employ the optimal-transport argument, as used by Gunsilius (2023), which estimates SC weights by minimizing the Wasserstein distance. However, in this study, we provide a simpler algorithm for estimating SC weights since the method of Gunsilius (2023) requires a procedure that is more complex than the original SCMs.
Because the mixture model (3) holds, we establish the following equation:
$\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\quad\forall t\in\mathcal{T},\ \ \forall\xi\in\mathbb{R},$ (4)
where $\sqrt{-1}$ denotes the imaginary number. From this equation, we estimate $w^{*}$ using the following methodology. We consider the upper bound of the difference between the expected exponential transforms of $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{j,t}$:
$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\right|\mathrm{d}\xi$
$=\int_{\xi\in(-h,h)}\left|\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right)+\epsilon|\xi^{G}|\right|\mathrm{d}\xi$
$\leq\int_{\xi\in(-h,h)}\left(\sum_{\gamma\in\{1,2,\dots,G\}}\frac{|\sqrt{-1}^{\gamma}||\xi|^{\gamma}}{\gamma!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\epsilon|\xi^{G}|\right)\mathrm{d}\xi$
$=\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}.$
This formula implies that by minimizing $\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}$, we can accurately estimate $\bm{w}^{*}$ such that $\int_{\xi in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\right|\mathrm{d}\xi=0$. This estimation strategy can be interpreted as a variant of the method of moments, where we estimate SC weights to match all moments between the treated unit and the weighted sum of the untreated units.
5.2 Core Assumptions
We make the following core assumption for identification of the SC weights.
Assumption 5.1.
The outcomes $(Y^{N}_{j,t})_{j\in\mathcal{J},t\in\mathcal{T}}$ are bounded.
Assumption 5.2.
There exists $\widetilde{\Phi}^{*}\in\Delta^{J}$ such that for all $\bm{w}^{*}\in\widetilde{\Phi}^{*}$, it holds that
$F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\quad\forall y\in\mathbb{R}.$
Assumption 5.1 allows us to guarantee the existence of the $G$-th moment of $Y^{N}_{j,t}$ for any $G\in\mathbb{N}$. Assumption 5.2 also has been utilized in existing studies such as Wan et al. (2018), Gunsilius (2023) and Shi et al. (2022). Wan et al. (2018) points out that the SCMs are based on Assumption 5.2. Gunsilius (2023) assumes Assumption 5.2 and estimates $\bm{w}^{*}$ by using minimizing the Wasserstein distance. Shi et al. (2022) investigates the underlying assumptions in the classical SCMs and then find that the combination of fine-grained models and Assumption 5.2 justifies the standard SCMs. We introduce the fine-grained models proposed by Shi et al. (2022) in Section 8.1.
Although the mixture model and linear factor models are different, they are closely related, as pointed out by Shi et al. (2022) and Nazaret et al. (2023). In terms of fine-grained models, Shi et al. (2022) identifies implicit assumptions underlying linear factor models, and we argue that the result justifies the use of a mixture model in the context of linear factor models. Furthermore, Nazaret et al. (2023) independently report this finding. For further details, see Section 8.1.
5.3 Estimation of SC Weights
Under the assumption of mixture models, we estimate the SC weight to match the characteristic functions (CHFs). We refer to our method as the Density Matching SCM (DMSCM).
Let $h\in(0,1)$ be a parameter determined and fixed by us. Then, we estimate SC weights as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|.$ (5)
By using the estimate $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}$, we can estimate the treatment effect and distributional treatment effect.
Treatment effect estimation.
Once estimating the weights, as well as the standard SCMs, we can estimate the treatment effect as
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}Y_{j,t}.$ (6)
We refer to this estimator as the Density Matching SCM (DMSCM) estimator. In Algorithm 1, we summarize our proposed DMSCM estimator with the D2MSCM estimator introduced in the following section.
Distributional treatment effect estimation.
We can also estimate the distributional treatment effect as the counterfactual distribution of $p_{0,t}(y)$ by a bootstrap method. We re-sample $L>1$ observations $\{Y^{l,\dagger}_{0}\}^{L}_{l=1}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$, (i) for each $j\in\mathcal{J}$, we re-sample $Y^{l,\dagger}_{j}$ from $\{Y_{j,t}\}_{t\in\mathcal{T}_{1}}$; (ii) we choose $Y^{l,\dagger}_{0}$ from $(Y^{l,\dagger}_{1},\dots,Y^{l,\dagger}_{J})$ with probability $(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}})_{j\in\mathcal{J}^{U}}$. In this case, our focus is on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$.
Algorithm 1 DMSCM.
  Parameter: The number of moments $G$.
  Obtain trained weights $\widehat{\bm{w}}^{\mathrm{DMSCM}}$ as (5).
  Estimate the treatment effect as (6).
6 Convergence Analysis
We provide the consistency results about our proposed SC weight estimator. The following theorem describes the results including the consistency for the treatment effect $\tau_{0}$.
Assumption 6.1.
Assume that the density function of $\varepsilon_{j,t}$ exists and is bounded for each $j\in\mathcal{J}$ and $t\in\mathcal{T}_{0}$, and that the stochastic process $\{\varepsilon_{j,t}\}^{T_{0}}_{t=1}$ is stationary, strongly mixing, with sum of mixing coefficient bounded by $M$.
Theorem 6.2.
Suppose that Assumptions 5.15.2 and 6.1 hold. Fix $h\in(0,1)$. Define
$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|$
Then, it holds that $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ as $T_{0}\to\infty$. Additionally, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$ and
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right),$
as $T_{0}\to\infty$ and $G\to\infty$.
The proof is shown in Appendix A. Note that this result can be applied for $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ as a special case. Thus, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ and $\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}$ are asymptotically unbiased under each assumption.
7 Inference
This section presents methods for statistical inference using our proposed DMSCM estimator.
7.1 Conformal Inference on the Treatment Effect
We consider statistical inference on the treatment effect using our proposed DMSCM and D2SCM to test a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$. Here, we focus on the conformal inference of Chernozhukov et al. (2021), which has been utilized in existing studies, such as Ben-Michael et al. (2021) and Ferman and Pinto (2021).
Our goal is to establish a confidence interval whose $p$-value $\widehat{p}$ satisfies that for any $a\in(0,1)$,
$\mathrm{Pr}\left(\widehat{p}\leq a\right)\to a$ (7)
as $T_{0}\to\infty$ and $T_{1}$ is fixed. To achieve this goal, the conformal inference takes the following three steps. First, for the sharp null hypothesis
$H_{0}:\tau_{0,t}=\alpha_{t},$
we create adjusted post-treatment outcomes for the treated unit $Y^{N\sharp}_{0,t}=Y^{I}_{0,t}-\alpha_{t}$ and add it to the original dataset as $\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$. Then, we apply our proposed method to the extended datasets consisting of
$\left(Y_{0,1},Y_{0,2},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$
and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$ to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and a corresponding predictor of the counterfactual outcome, denoted by $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$. Finally, we compute the $p$-value by assessing whether the adjusted residual “conforms” with the pretreatment residuals. Let $u_{t}(\bm{\alpha})$ be the residual defined as $u_{t}(\bm{\alpha})=Y_{0,t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{0}$ and $u_{t}(\bm{\alpha})=Y^{I}_{0,t}-\alpha_{t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{1}$. For $j\in\mathcal{T}$, let $\pi_{j}$ be a permutation of indexes defined as $\pi_{j}(t)=t+j$ if $t+j\leq T$ and $\pi_{j}(t)=t+j-T$ otherwise. Let $\Pi=\{\pi_{1},\dots,\pi_{T-1}\}$. Then, the $p$-value is defined as $\widehat{p}(\bm{\alpha})=1-\widehat{F}(\bm{\alpha})$, where
$\widehat{F}(\bm{\alpha})=\frac{1}{T}\sum_{j\in\mathcal{T}}\mathbbm{1}\left[\frac{1}{T}\sum_{t\in\mathcal{T}}\left|u_{t}(\bm{\alpha})\right|<\frac{1}{T}\sum_{\pi_{j}\in\mathcal{T}}\left|u_{\pi_{j}}(\bm{\alpha})\right|\right].$
Because $Y^{N}_{0,t}$ is random, by inverting this test, we can construct a confidence interval for $\tau_{0,t}$, which is equivalent to constructing a conformal prediction set (Vovk et al., 2005) for $Y^{N}_{0,T}$. The $(1-\xi)$ confidence interval is given as
$\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\mid\widehat{p}(\bm{\alpha})>\xi\right\},$
where $\mathcal{A}$ is a set of candidates of $\bm{a}$. Theorem 1 of Chernozhukov et al. (2021) guarantees (7).
7.2 Two-Sample Homogeneity Test
We define the null and alternative hypotheses as $H_{0}:p^{I}_{t}=p^{N}_{t}$ and $H_{1}:p^{I}_{t}\neq p^{N}_{t}$. Then, we can conduct hypothesis testing by applying two-sample homogeneity test, and our distributional treatment effect estimator. For example, we can employ maximum mean discrepancy (MMD) for testing (Gretton et al., 2012).
8 Discussion
This section discusses some remaining problems.
8.1 Justification of Mixture Models
To elucidates assumptions in SCMs, Shi et al. (2022) posits fine-grained models, where each unit has its components, and the outcomes of treated and untreated units are averages of the outcomes of the components. For example, in the empirical study of the tobacco control program in California, the outcome of interest is annual per capita cigarette consumption at the state level, which is an average of the consumption of people in the state. Based on their arguments, we can justify the assumption of mixture models by connecting them to linear factor models.
Assume that there are $U_{j}$ components in unit $j\in\mathcal{J}$. For each $u\in\mathcal{U}_{j}\coloneqq\{1,2,\dots,U_{j}\}$, let $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ an outcome of a component $u$ in unit $j$ at time $t$. We also suppose an existence of a $d_{W}$-dimensional unobserved random variable $W_{j,t,u}\in\mathbb{Z}^{d_{W}}$, contributing to individual $u$’s outcome, such as age, education level, or income. We refer to $W_{j,t,u}$ as causes. Assume that $(Y^{I}_{j,t,u},Y^{N}_{j,t,u},W_{j,t,u})$ is an i.i.d. random variable across $u\in\mathcal{U}_{j}$.
The relationship between the causes $W_{j,t,u}$ and potential outcomes $(Y^{I}_{j,t,u},Y^{U}_{j,t})$ can be linear or non-linear, and can also be time-variant. Let $p_{j,t}(w,y)$ be the PDF of $(W_{j,t},Y^{N}_{j,t})$, $p_{j,t}(y\mid w)$ be the PDF of $Y^{N}_{j,t}$ given $W_{j,t}$, and $p_{j,t}$ be the marginalized PDF of $W_{j,t}$. The authors assume that interventions are made at a group-level and individuals in each group comply with their group-level intervention.
Under this setting, Shi et al. (2022) finds that the following assumptions are required for the linearity of SCMs (2).
Assumption 8.1 (Independent Causal Mechanism; ICM).
Conditional on the causes $W_{j,t,u}$, the potential outcome $Y^{N}_{j,t,u}$ is independent of the unit $j\in\mathcal{J}$. For population distribution $j\in\in\mathcal{J}$ at time $t\leq\mathcal{T}$, the joint distribution of $(Y^{N}_{j,t,u},W_{j,t,u})$ is given as $p_{j,t}(w,y)=r_{j,t}(w)p_{t}(y|w)$.
Assumption 8.2 (Stable distributions).
Decompose the causes into two subsets $W_{j,t,u}=\{U_{j,t,u},S_{j,t,u}\}$. Let $S$ denote the subset that differentiates the treated unit from the selected untreated units (minimal invariant set); that is, its distribution in the target unit is different from its distribution in the selected untreated units. We assume that for all groups, the distribution of $S$ does not change for all time periods $t\leq T$, $r_{j,t}(w)=\zeta_{j}(s)\xi_{t}(u\mid s)$, where $\zeta_{j}(s)$ is the PDF of $S_{j,t,u}$, and $\xi_{t}(u\mid s)$ is the PDF of $U_{j,t,u}$ given $S_{j,t,u}$.
Assumption 8.3 (Sufficiently similar untreated units).
Let $S$ be the minimal invariant set for the target and untreated units. The untreated units are sufficiently similar if the cardinality of the minima invariant set, $\big{|}\mathcal{J}^{U}\big{|}\geq|S|$.
Assumption 8.4 (Common support).
Let $(s_{1},\dots,s_{R})$ be the support of $S$. For each $s$, there exists at least one untreated unit $j$, where $\zeta_{j}(S=s)>0$.
Proposition 8.5 (Causal identifiability. From Theorem 1 of Shi et al. (2022)).
Suppose that Assumptions 8.18.4 hold. Then, there exists $\{w^{*}_{j}\}_{j\in[J]}$ such that $p^{N}_{0,t}(y)=\sum_{j\in\mathcal{U}^{U}}w^{*}_{j}p^{N}_{j,t}(y)$ holds.
This proposition also implies that the mixture model holds (Nazaret et al., 2023), which is consistent to the linear factor models. Therefore, based on this result, we can obtain the following result for linear factor models as well as Ferman and Pinto (2021).
Corollary 8.6.
Suppose that Assumptions 5.1, 6.1, and 8.18.4 hold. Then, $\bm{w}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}$ holds as $T_{0}\to\infty$, and $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$. Additionally, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ converge to $\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)$ in probability as $T_{0}\to\infty$ and $G\to\infty$.
The proof is straightforward. From Proposition 8.5, under Assumptions 8.18.4, (3) holds. Therefore, from Theorem 6.2, the claim holds under (1) with $c_{j}=0$ for $j\in\mathcal{J}$.
8.2 Auxiliary Covariates
We can consider a case where auxiliary covariates are observable. For each unit $j\in\mathcal{J}$, let
$X_{j,t}\coloneqq(X_{j,t,1},X_{j,t,1},\dots,X_{j,t,K})\in\mathbb{R}^{K}$
be $K$-dimensional covariates, which is not affected by treatment for all $j\in\mathcal{J}$. Let
^$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left\{X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right\}.$
When the auxiliary covariates are observable, we extend the DMSCM estimator as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}\coloneqq v_{0}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|+\sum^{K}_{k=1}v_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
where $\bm{v}\coloneqq(v_{k})_{k\in\{0,1,2,\dots,K\}}$ are some weights such that $\sum^{K}_{k=0}v_{0}=1$.
8.3 Distributional SCMs and Optimal Transports
Gunsilius (2023) establies the framework of distributional SCMs. In the framework, that study assumes the following:
$F^{-1}_{Y_{0,t},N}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y_{j,t},N}(q),\ \ \ \forall q\in(0,1),$
where $F^{-1}_{Y_{j,t},N}(q)$ is the quantile function defined as $F^{-1}_{Y_{j,t},N}(q)\coloneqq\inf_{y\in\mathbb{R}}\big{\{}F_{Y_{j,t},N}(q)\geq q\big{\}}$. Then, Gunsilius (2023) estimates the SC weight to match the distribution under the Wasserstein distance. Their method is called DiSCo. Specifically, they estimate the SC weights as
$\widehat{\bm{w}}^{\mathrm{DiSCo}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\frac{1}{M}\sum^{M}_{m=1}\left|\widehat{F}^{-1}_{Y^{N}_{0,t}}(V_{m})-\sum_{j\in\mathcal{J}^{U}}w_{j}\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})\right|,$
where $\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})$ is an empirical quantile function of $Y^{N}_{j,t}$, and $\{V_{m}\}$ is a set of $M$ independent draws from the uniform distribution with support $[0,1]$.
8.4 Demeaned Density Matching SCM (D2MSCM)
Next, we consider a case where the target distribution does not follow mixture models, but the deviation is zero when it is marginalized. Instead of (3), We consider the following another mixture model
$p^{N}_{0,t}(y-c_{0})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y-c_{j})\qquad\forall t\in\mathcal{T}.$
From the viewpoints of linear factor models 1 and fine-grained models (Shi et al., 2022), this assumption allows us $c_{j}\neq 0$ for some $j\in\mathcal{J}$. Under the previous mixture model, we only consider a case where $c_{j}=0$ for all $j\in\mathcal{J}$, which is restrictive in practice.
To estimate the SC weight, we first estimate $c_{j}$ by taking the average of $(Y_{j,t})_{t\in\mathcal{T}_{0}}$ and subtract it from $Y_{j,t}$. This manipulation corresponds to the demeaned estimation in Ferman and Pinto (2021). Let $h\in(0,1)$ be a parameter determined and fixed by us. Then, we estimate SC weights as
$\widehat{\bm{w}}^{\mathrm{D2MSCM}}_{T_{0}}\coloneqq\left(\widehat{w}^{\mathrm{D2MSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{0,t}-\overline{Y}^{N}_{0}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}-\overline{Y}^{N}_{0}\Big{)}^{\gamma}\right)\right|.$
We refer to this method as the demeaned DMSCM (D2SCM). By using the D2MSCM estimator $(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}})_{j\in\mathcal{J}^{U}}$, we estimate the treatment effect as
$\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\left(\overline{Y}^{N}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{D2SCM}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\right).$ (8)
9 Experiments
We conduct experiments with simulated and real-world data. Additional results are in Appendix 9.1.
9.1 Simulation Studies
In this section, we perform simulation studies to show how our proposed SCMs compared to traditional SCMs. Let $J\in\{10,30,60\}$ and $G\in\{2,3,5,10,20,30,40,50,60,70,80,90,100\}$, $T_{0}=30$, $T_{1}=100$, and $K=5$. We perform a simulation on SCMs for $p_{0,t}(\bm{z})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p_{0,t}(\bm{z})$, where $\bm{z}$ is ($K+1$)-dimensional ($6$-dimensional) vector, where the first element corresponds to $Y_{j,t}$ and the rest corresponds to $X_{j,t}$. For each $j\in\mathcal{J}$, let $p_{j,t}$ be a density of a multinomial normal distribution $\mathcal{N}(\bm{\mu}_{j,t},\bm{\sigma}^{2}_{j,t})$, where $\bm{\mu}_{j,t}=(\mu_{j,t,1},\mu_{j,t,2},\mu_{j,t,3},\mu_{j,t,4},\mu_{j,t,5},\mu_{j,t,6})$, and $\bm{\sigma}^{2}_{j,t}$ is a $(6\times 6)$-diagonal matrix with a $(k,k)$-element denoted by $\sigma^{2}_{j,t,k}$. Each mean parameter is drawn from a standard normal; distribution. Each variance parameter is drawn from a uniform distribution with a support $[1,20]$. Let us assume that for each $j\in\mathcal{J}$, $\mu_{j,t,1}=\mu_{j,t,1}+\epsilon_{j,t}$ and $\sigma^{2}_{j,t,1}=\sigma^{2}_{j,t,1}+\widetilde{\epsilon}_{j,t}$, where $\epsilon_{j,t}$ and $\widetilde{\epsilon}_{j,t}$ follow a normal densities $\mathcal{N}(0,10)$, respectively. Let us correct the variance as $\widetilde{\epsilon}_{j,t}=0.1$ if $\widetilde{\epsilon}_{j,t}\leq 0.1$ occurs. For the treated unit, we assume that $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$; that is $\tau_{0,t}=20$. We draw $\bm{w}^{*}$ from a uniform distribution and normalize the sum to be $1$. We apply DMSCM and Abadie $1,000$ times and compute the error of the treatment effect estimation. We show the result with boxplot in Figure 4. Our proposed method successfully estimate the treatment effect with smaller estimation error. The estimation decreases as $G$ increases.
We conduct simulation studies to illustrate the comparative performance of our proposed SCMs and traditional SCMs within a setting analogous to Section 9.1. The only modifications introduced involve the parameters. Specifically, $J\in\{1,5,10,15,20,\dots,35,40,45,50\}$, $G\in\{2,3,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$. All other conditions align with those detailed in Section 9.1.
We apply DMSCM and Abadie $1,000$ times and compute the error of the treatment effect estimation in Figure 7. Our focus lies on the fluctuation in mean squared estimation errors as the number of units, $J$, increases. Despite the method by Abadie (2002) incurring escalating estimation errors with the increase of $J$, estimation errors do not exhibit a similar growth under our proposed methods. Furthermore, our proposed methods exhibit lower estimation errors compared to those under the method of Abadie (2002).
A comparable phenomenon is evident in distributional treatment effect estimation. We compare the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability distance computed through kernel embedding. The upper figure in Figure 8 displays how MMDs alter as $J$ increases. As illustrated, our proposed methods outperform even in this task. The lower figure in Figure 8 presents results for $G\in\{2,10,50,100\}$. As the results indicate, the predictive capability augments as $G$ (the number of moment conditions) increases.
Our findings might have a relationship with those of Spiess et al. (2023), who investigates high-dimensional SCMs. Future investigations will be devoted to exploring these connections.
Figure 2: Results of simulation studies.
Figure 3: Counterfactual outcomes in empirical analysis.
Figure 4: Counterfactual distributions in empirical anaylsis.
Table 2: Results of conformal inference. Estimated treatment effect $\widehat{\tau}_{0,t}$ and $90\%$ confidence interval $(\underline{\tau}_{0,t},\overline{\tau}_{0,t})$.
Terrorist conflict in Basque country Tobacco control program in California Reunification of Germany
Abadie DMSCM Abadie DMSCM Abadie DMSCM
$\gamma$ $2$ $5$ $10$ $100$ $2$ $5$ $10$ $100$ $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$0.800 -0.845 -1.054 -0.279 -0.138 -26.148 -27.744 -32.306 -35.080 -36.879 -1492.374 -1253.616 -1419.120 -1324.448 -510.765
$\overline{\tau}_{0,t}$ 0.522 0.699 0.495 0.521 0.578 8.698 7.852 8.898 9.544 11.748 2635.583 -1008.822 -591.420 -716.376 2327.944
$\underline{\tau}_{0,t}$2.587 -2.150 -2.605 -2.616 -2.590 -36.798 -29.268 -34.783 -37.308 -35.243 -2390.413 -2738.232 -1989.323 -2018.878 -2625.129
Figure 5: Results of simulation studies on ATE prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the mean squared prediction error.
Figure 6: Results of simulation studies on distributional treatment effect prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the MMD.
Figure 7: Results of simulation studies on ATE prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the mean squared prediction error.
Figure 8: Results of simulation studies on distributional treatment effect prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the MMD.
9.2 Empirical Analysis
We conduct empirical studies discussed in Abadie and Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015).
Terrorist conflict in the Basque Country. Abadie and Gardeazabal (2003) investigates the effect of terrorist conflict in the Basque Country on gross domestic product (GDP). In this example, $J=16$, $K=14$, and $(t_{0},T_{0},T_{1})=(1955,1970,1997)$. Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain. ETA promoted an independent Basque state by carrying out terrorist activities mainly in the Basque Country region of Spain. Abadie and Gardeazabal (2003) utilizes annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA conducted significant terrorist activity. In the data, 15 years of pre-intervention data are included. The set of untreated units comprises $17$ other Spanish regions. Each region’s datum contains the outcome of interest (GDP per capita) at the regional level, with covariates including investment rate (percentage), population density, five sectoral productions, and four human capitals, all averaged over the period from 1955 to 1968.
Tobacco control program in California. Abadie et al. (2010) study the effect of a large tobacco control program adopted in California in $1988$, which we explained in our introduction. In this example, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$. Abadie et al. (2015) investigates the treatment effect of the reunification of Germany. East Germany was absorbed into West Germany, the result being a reunited Germany in 1990. They utilize annual panel data at the country level from 1960 to 2003. The outcome of interest is GDP per capita at the country level. The pre-intervention period spans $30$ years. The set of untreated units comprises $16$ OECD countries that make up synthetic West Germany. Covariates include inflation rate, industry share, investment rate, schooling, and a measure of trade openness.
Reunification of Germany. Abadie et al. (2015) studies the effect of the reunification of Germany on GDP in 1990. Because the countries were dissimilar from any one country to make a comparison group, they apply SCMs to create a composite comparison group. In this example, $J=8$, $K=31$, and $(t_{0},T_{0},T_{1})=(1960,1990,2003)$. Abadie et al. (2010) investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California. They utilize annual state-level panel data from 1970 to 2000. The program was implemented in January 1989, providing 18 years of pre-intervention data. Four states that have adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that increased state cigarette taxes by $50$ cents or more between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) and the District of Columbia are excluded from the untreated units of 38 states. The outcome of interest is annual per capita cigarette consumption at the state level, measured as per capita cigarette sales in packs. Covariates include the average retail price of cigarettes, state income per capita (log), percentage of the population aged 15-24, and beer consumption per capita; all averaged over the period from 1980 to 1988.
Results. We show the result of the estimated treatment effect and $90$% confidence interval in Table 2. We show the results of treatment effect estimation in Figure 4 and distributional treatment effect estimation in Figure 4.
For each dataset, we plot the counterfactual outcomes and distributions in Figures 4 and 4, respectively. We also conduct conformal inference on the sharp null hypothesis $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect. For the results of Abadie, DMSCM, We show the $90\%$ confidence interval of the treatment effect in Table 2. Our proposed method has a good fit in terms of the observed outcomes in pretreatment periods $t\in\mathcal{T}_{0}$. Therefore, we consider that our proposed method also predicts counterfactual outcomes and distributions well. In conformal inference, our method does not reject the null hypothesis as with Abadie, except for the DMSCM in the tobacco control problem in California.
10 Conclusion
In this study, we identified the problem of implicit endogeneity in the linear regression models commonly used in existing SCM studies. To address this issue, we proposed a novel approach to SCMs using density matching. Our proposed method enables consistent estimation of SC weights and improves the accuracy of counterfactual outcome predictions compared to existing methods. We validated the effectiveness of our approach through simulation experiments and empirical analyses.
Appendix A Proof of Theorem 6.2
First, we confirm the basic fact about the uniqueness of the CHF.
Proposition A.1 (Theorem 26.2 in Billingsley (1995)).
Distinct measures cannot have the same characteristic function.
Let us define
$\widehat{Q}_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|,$
$Q_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|.$
Proof.
First, from the definition of $\Delta^{J}$, $\Delta^{J}$ is a compact parameter space. Next, $\widehat{Q}(\bm{w})$ converges uniformly to $Q(\bm{w})$ as $T_{0}\to\infty$ under Assumption 6.1. Furthermore, $Q(\bm{w})$ is continuous with respect to $\bm{w}$.
Under these conditions, for any $\varepsilon>0$, we have with probability approaching one, we have
  1. 1.
    $\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}(\bm{w}_{0})-\varepsilon/3$;
  2. 2.
    $\widehat{Q}_{G}(\widehat{\bm{w}}^{\mathrm{DMSCM}})>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3$;
  3. 3.
    $\widehat{Q}_{G}\left(\bm{w}_{0}\right)>Q_{G}(\bm{w}_{0})-\varepsilon/3$.
Therefore, with probability approaching one,
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3>\widehat{Q}_{G}\left(\bm{w}_{0}\right)-2\varepsilon/3>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon.$
Thus, for any $\varepsilon>0$, $Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon$ with probability approaching one. Let $\mathcal{N}$ be any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Since $\Theta\cap\mathcal{N}^{c}$ is compact, and $Q_{G}(\bm{w})$ is continuous with respect to $\bm{w}$, we have
$\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})=Q_{G}(\bm{w}^{\dagger})<Q_{G}(\bm{w}_{0})$
for some $\bm{w}^{\dagger}\in\Theta\cap\mathcal{N}^{c}$. Thus, choosing $\varepsilon=Q_{G}(\bm{w}_{0})-\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})$, it follows that
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w}),$
and hence $\widehat{\bm{w}}^{\mathrm{DMSCM}}\in\mathcal{N}$, where recall that $\mathcal{N}$ is any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Thus, $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ holds as $T_{0}\to\infty$. Additionally, from Proposition A.1, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$. ∎