### v4

# Asymptotically Unbiased Synthetic Control Methods by Density Matching

Abstract
Synthetic Control Methods (SCMs) have become a fundamental tool for comparative case studies. The core idea behind SCMs is to estimate treatment effects by predicting counterfactual outcomes for a treated unit using a weighted combination of observed outcomes from untreated units. The accuracy of these predictions is crucial for evaluating the treatment effect of a policy intervention. Subsequent research has therefore focused on estimating SC weights. In this study, we highlight a key endogeneity issue in existing SCMs—namely, the correlation between the outcomes of untreated units and the error term of the synthetic control, which leads to bias in both counterfactual outcome prediction and treatment effect estimation. To address this issue, we propose a novel SCM based on density matching, assuming that the outcome density of the treated unit can be approximated by a weighted mixture of the joint density of untreated units. Under this assumption, we estimate SC weights by matching the moments of the treated outcomes with the weighted sum of the moments of the untreated outcomes. Our method offers three advantages: first, under the mixture model assumption, our estimator is asymptotically unbiased; second, this asymptotic unbiasedness reduces the mean squared error in counterfactual predictions; and third, our method provides full densities of the treatment effect rather than just expected values, thereby broadening the applicability of SCMs. Finally, we present experimental results that demonstrate the effectiveness of our approach. [^1] This study was presented at the Workshop on Counterfactuals in Minds and Machines at the International Conference on Machine Learning in July 2023 and at the International Conference on Econometrics and Statistics in August 2023.
1 Introduction
Synthetic control methods (SCMs, Abadie and Gardeazabal, 2003; Abadie et al., 2010) have emerged as a crucial tool for program evaluation in comparative case studies across disciplines such as economics, statistics, and machine learning. As noted by Athey and Imbens (2017), SCMs represent “arguably the most important innovation in the policy evaluation literature in the last 15 years” due to their “simplicity of the idea, and the obvious improvement over standard methods.”
In SCMs, there are multiple units, where one unit receives a policy intervention (treated unit) at a certain time while the remaining units remain untreated. The objective is to estimate the treatment effect for the treated unit, defined as the difference between its observed factual outcome and its unobserved counterfactual outcome. Untreated units help account for unobserved trends in the outcome over time that are unrelated to the policy intervention. The key idea is that an optimally weighted combination of the untreated units, referred to as the synthetic control (SC) unit, serves as an appropriate estimate of the counterfactual outcome. See Figure 1 for an illustration. d SCMs have been widely applied in various empirical studies, such as evaluating the impact of terrorism on GDP (Abadie and Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham and Shah, 2017). For instance, Abadie et al. (2010) applies SCM to study the treatment effect of a comprehensive tobacco control program implemented in California in 1988. While smoking rates in California declined following the program’s implementation, it remained unclear whether this decrease was attributable to the policy. For estimating the treatment effect, annual per-capita cigarette sales data from states that did not implement a similar tobacco control program were used as untreated units. SCM was then applied to construct California’s counterfactual outcome, representing the smoking rate in the absence of the tobacco control program.
A key challenge in SCMs is the assumption, made in most existing studies (Abadie, 2002), that the expected outcome of the treated unit can be represented as a weighted sum of the outcomes of untreated units. Ferman and Pinto (2021) highlights that this assumption leads to a non-negligible bias in the estimator. According to Ferman and Pinto (2021), this bias arises from endogeneity in the linear regression model implied by the SCM assumption due to the correlation between the outcomes of untreated units and the error term. This issue is analogous to the endogeneity caused by measurement error (Greene, 2003). We refer to this problem as implicit endogeneity.
To address this issue, we propose an alternative approach based on a mixture model assumption rather than a linear relationship among expected outcomes. Specifically, we assume that the density of the treated unit’s outcome can be approximated by a weighted combination of the densities of the untreated units. This assumption further implies that each moment of the treated unit’s outcome can be represented as a weighted sum of the moments of the untreated units. Leveraging this insight, we estimate SC weights by matching the moments of the treated unit’s outcome to those of the untreated units. We refer to our method as the Density Matching SCM (DMSCM). Unlike existing methods, our proposed estimator is asymptotically unbiased if the mixture model assumption is correct. We validate the effectiveness of our approach through both simulation studies and empirical analysis.
The mixture model assumption has appeared in previous and concurrent studies, including Wan et al. (2018), Gunsilius (2023), and Shi et al. (2022). Our proposed method can be viewed as a simplification of the distributional SCMs introduced by Gunsilius (2023). While Gunsilius (2023) does not explicitly address the implicit endogeneity problem, they note that their results diverge from those obtained under standard SCMs. This discrepancy arises because, under the mixture model assumption, Gunsilius (2023) provides a consistent estimator for the treatment effect, whereas existing SCM methods do not. We establish a connection between distributional SCMs and the asymptotic bias problem and propose a simplified estimation method based on density matching.
In summary, our contributions are as follows:
  • Reexamining the assumptions necessary for treatment effect estimation in SCMs.
  • Proposing a density-matching approach for asymptotically unbiased SCMs.
  • Simplifying the implementation of density matching in SCMs.
We formulate the problem in Section 2. Next, in Section 3, we review the basic idea of SCMs. In Section 4, we highlight the problem of implicit endogeneity. In Section 5, we then propose our method under the mixture model assumption. We investigate the asymptotic properties of this method in Section 6 and present a statistical inference method in Section 7. In Section 8, we discuss related topics and existing studies. Finally, in Section 9, we examine the performance of our estimators through extensive simulations.
Figure 1: Illustration of SCMs
Related work. Following the pioneering work by Abadie and Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015). These SCMs, including their variants such as those by Doudchenko and Imbens (2016) and Ben-Michael et al. (2019), are based on minimizing the pre-treatment sum of squared errors between the outcome of the treated unit and the weighted sum of those of the untreated units, with some regularization and constraints regarding the weights. We refer to such SCMs as Least-Squares SCMs (LS-SCMs).
The bias of the LS-SCMs strongly depends on whether there exist weights that can predict the outcome of the treated unit without error. In other words, a random outcome of the treated unit must be replicable by the sum of outcomes of untreated units. Since such weights yield a perfect fit in the estimation process using observations from the pre-treatment period, this situation is called pretreatment-fit.
Abadie et al. (2010) demonstrates that under a pre-treatment fit, the bias of the LS-SCM estimator converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases. However, Ferman and Pinto (2021) finds that if the pre-treatment fit does not hold, then the LS-SCM estimator has a bias and is not consistent. The cause of this bias is the correlation between the outcomes of the untreated units and the error term in the posited linear model, which we refer to as implicit endogeneity.
Several methods have been proposed to address this bias issue. Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where both the number of untreated units ($J\to\infty$) and the length of the pre-intervention periods ($T_{0}$) grow large. Concurrently with us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias reported by Ferman and Pinto (2021). [^2] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and he and we conducted our studies completely independently. While our formulation and his regarding endogeneity are similar, our approach differs significantly. We address the problem through density matching, similar to the strategy in Gunsilius (2023), whereas Fry (2024) employs an instrumental variable approach using outcomes of untreated units not used in SCMs as instruments.
In this study, we focus on mixture models among outcome distributions, an assumption also employed in Wan et al. (2018), Gunsilius (2023), Shi et al. (2022), and Nazaret et al. (2023). Although these studies introduce mixture models with motivations other than the bias problem, we highlight that such models can effectively circumvent bias under imperfect pre-treatment fit.
Here, we briefly review existing work using mixture models. Wan et al. (2018) points out that SCMs essentially rely on mixture models, distinguishing them from traditional panel data analysis. Gunsilius (2023) proposes distribution SCMs (DiSCo), which estimate SC weights by matching the outcome distributions of the treated unit and the weighted sum of the distributions of the untreated units. Shi et al. (2022) introduces fine-grained models to elucidate hidden assumptions in SCMs and explores connections to linear factor models. Nazaret et al. (2023) extends these insights by analyzing model misspecification under a mixture model formulation.
This study also addresses a question raised by Gunsilius (2023). While Gunsilius (2023) suggests that standard least squares SCMs try to match the first moment of the outcome distribution (and therefore differ from distributional approaches), we stress that least squares approaches are additionally vulnerable to implicit endogeneity, which can generate bias. Our density-matching approach avoids this problem by matching a broader set of distributional features.
Table 1: Comparison of SCMs. See the following sections for the definitions of variables and methods.
Method Assumption Asymptotic unbiasedness Remark
Ours DMSCM $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\ \ \ \forall y\in\mathbb{R}$ Unbiased
Abadie and Gardeazabal (2003) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Biased
Ferman (2021) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Unbiased $J\to\infty$
Gunsilius (2023) DiSCo $F^{-1}_{Y^{N}_{0,t}}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y^{N}_{j,t}}(q)\ \ \ \forall q\in(0,1)$ Unbiased
Fry (2024) IV method (GMM) $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Unbiased Use IVs
2 Problem Setting
Suppose there are $J+1$ units, indexed by $j\in\mathcal{J}\coloneqq\{0,1,2,\dots,J\}$, and a time series indexed by $t\in\mathcal{T}=\{1,2,\dots,T\}$, where $3\leq T\in\mathbb{N}$. Without loss of generality, assume that the first unit ($j=0$) is the treated unit, i.e., the unit affected by the policy intervention of interest. The set of potential control units is $\mathcal{J}^{U}\coloneqq\mathcal{J}\backslash\{0\}=\{1,2,\dots,J\}$, consisting of units not affected by the intervention. We also assume that our data span $T$ periods and that the intervention occurs at $t=T_{0}$, with $2\leq T_{0}<T$. Observations in period $t\in\mathcal{T}_{0}\coloneqq\{1,2,\dots,T_{0}\}$ are from before the intervention, and observations in period $t\in\mathcal{T}_{1}\coloneqq\mathcal{T}\backslash\mathcal{T}_{0}$ are from after the intervention. Let $T_{1}\coloneqq|\mathcal{T}_{1}|=T-T_{0}$.
2.1 Potential Outcomes
Following the Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974), we define the potential outcomes of the policy intervention. For each unit $j\in\mathcal{J}$ and period $t\in\mathcal{T}$, there are two outcomes of interest, $\left(Y^{I}_{j,t},Y^{N}_{j,t}\right)\in\mathbb{R}^{2}$, corresponding to scenarios with and without the intervention. For each $j\in\mathcal{J}$ and $t\in\mathcal{T}$, let $F^{I}_{j,t}(y)$ and $F^{N}_{j,t}(y)$ be the cumulative distribution functions (CDFs) for $Y^{I}_{j,t}$ and $Y^{N}_{j,t}$, respectively.
2.2 Observations
We next define our observations. For each unit $j\in\mathcal{J}$, only one of the two potential outcomes can be observed, corresponding to what actually happened. Specifically, for each $j\in\mathcal{J}$, we observe $Y_{j,t}\in\mathbb{R}$ such that
$Y_{0,t}=\begin{cases}Y^{N}_{0,t}&\forall t\in\mathcal{T}_{0},\\ Y^{I}_{0,t}&\forall t\in\mathcal{T}_{1},\end{cases}\quad\mathrm{and}\quad Y_{j,t}=Y^{N}_{j,t}\quad\forall j\in\mathcal{J}^{U},\forall t\in\mathcal{T}.$
We aim to estimate the effect of the policy intervention by predicting $Y^{I}_{0,t}$ from $\left(Y^{N}_{j,t}\right)_{j\in\mathcal{J}^{U}}$ and comparing that prediction to the observed $Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$. We define the treatment effect more formally in the next subsection.
2.3 Treatment Effect
In SCMs, there are several ways to define the treatment effect $\tau_{0,t}\in\mathbb{R}$ for each $t\in\mathcal{T}_{1}$ (i.e., after the intervention). First, for each $t\in\mathcal{T}_{1}$, we can define the treatment effect for the treated unit $j=0$ as
$\tau_{0,t}\coloneqq\mathbb{E}[Y^{I}_{0,t}]-\mathbb{E}[Y^{N}_{0,t}].$
Another definition is based on the following linear factor model:
Definition 2.1.
We say that $(Y^{I}_{j,t},Y^{N}_{j,t})$ follows the linear factor model if:
$Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t},\quad\mathbb{E}_{j,t}[\epsilon_{j,t}]=0,\quad\forall t\in\mathcal{T},\quad\forall j\in\mathcal{J},$ (1)
$Y^{I}_{j,t}=\tau_{j,t}+Y^{N}_{j,t},$
where $\delta_{t}$ is an unobserved common factor with constant factor loadings across units, $c_{j}$ is an unknown time-invariant fixed effect, $\lambda_{t}$ is a $(1\times F)$ vector of unobserved common factors, $\mu_{j}$ is a $(F\times 1)$ vector of unknown factor loadings, and $\epsilon_{j,t}$ is an unobserved idiosyncratic shock. Here, $c_{j}$, $\mu_{j}$, and $\tau_{0,t}$ are non-random variables, while $\delta_{t}$, $\lambda_{t}$, and $\epsilon_{j,t}$ are random variables.
Under the linear factor model, the treatment effect is given by $\tau_{0,t}=Y^{I}_{0,t}-Y^{N}_{0,t}$, a non-random variable by definition. Consequently, the expression $\tau_{0,t}\coloneqq\mathbb{E}[Y^{I}_{0,t}]-\mathbb{E}[Y^{N}_{0,t}]$ also holds under the factor model.
Comparative case studies aim to evaluate $\tau_{0,t}$ by predicting $Y^{N}_{0,t}$ for $t>T_{0}$, the counterfactual outcome that would have been observed for the treated unit had the intervention not occurred. Because the intervention takes place after $t=T_{0}$, $Y^{N}_{0,t}$ is a counterfactual outcome. In settings where data consist of only a few units (e.g., regions or countries), it can be difficult to find a single untreated unit that provides a suitable comparison. For this problem, SCMs enable us to construct a counterfactual outcome using a predictor based on the outcomes of the untreated units. Under suitable model assumptions that relate $Y^{N}_{0,t}$ to $\left(Y^{N}_{j,t}\right)_{j\in\mathcal{J}^{U}}$, we can construct the predictor using observations from the pre-intervention period and then predict the unobserved counterfactual outcome in the post-intervention period. Consequently, we estimate the treatment effect by comparing the predicted $Y^{N}_{0,t}$ with the observed $Y^{I}_{0,t}$.
Distributional Treatment Effects. In program evaluation, we are often interested not only in the scalar treatment effect but also in the distributions of outcomes (Maier, 2011) or their functionals, such as quantile treatment effects (QTEs) and expected social welfare (Abadie, 2002). By applying our method, we can also estimate the distribution of the counterfactual outcome, $p^{N}_{0,t}(y)$, for $t\in\mathcal{T}_{1}$. The treatment effect is then captured by the difference between $p^{I}_{0,t}(y)$ (the distribution of the actual outcome) and $p^{N}_{0,t}(y)$, or by functionals of these distributions, such as QTEs and expected social welfare. Depending on the application, we may also adjust the target density, for example, by averaging across all post-intervention periods: $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$. We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus and Oprescu, 2022; Chikahara et al., 2022). In the context of SCMs, Gunsilius (2023) and Chen (2020) have proposed methods for estimating such distributional treatment effects.
3 Recap of SCMs
This section reviews the original SCM proposed by Abadie and Gardeazabal (2003). To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ using a weighted sum of $\{Y^{N}_{j,t}\}_{j=1}^{J}$ for $t\in\mathcal{T}_{1}$. Specifically, SCMs predict $Y^{N}_{0,t}$ as
$\widehat{Y}^{N}_{0,t}\coloneqq w_{0}+\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t},$
where $w_{0}\in\mathbb{R}$ and $w_{j}\in\mathbb{R}$ $(j\in\mathcal{J}^{U})$ are the SC weights.
Several methods exist for estimating the SC weights. In the SCM proposed by Abadie and Gardeazabal (2003), these weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. Define
$\Delta^{J}\coloneqq\left\{\bm{w}=(w_{1}\ w_{2}\ \dots\ w_{J})^{\top}\in[0,1]^{J}\mid\sum_{j\in\mathcal{J}^{U}}w_{j}=1\right\}.$
Then, Abadie and Gardeazabal (2003) estimate the SC weights as
$\bm{w}^{\mathrm{Abadie}}=\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t}\right)^{2}.$
Recall that we referred to SCMs using this approach as the Least-Squares-SCMs (LS-SCMs). The counterfactual outcome is then predicted as
$\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}=\widehat{w}^{\mathrm{Abadie}}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{Abadie}}_{j}Y^{N}_{j,t}.$
Using $\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$, we estimate the treatment effect as
$\widehat{\tau}^{\mathrm{Abadie}}_{t}=Y^{I}_{0,t}-\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}.$
The literature on inference for SCMs has grown significantly, extending the original proposals by Abadie and Gardeazabal (2003) and Abadie et al. (2010). Notable contributions include works by Li (2020), Shaikh and Toulis (2021), Cattaneo et al. (2021), and Chernozhukov et al. (2019). Doudchenko and Imbens (2016) proposes estimating the SC weight as
$\widetilde{\bm{w}}^{\mathrm{DI}}=\operatorname*{arg\,min}_{\widetilde{\bm{w}}=(w_{0}\ w_{1}\ \cdots\ w_{J})^{\top}\in\mathbb{R}^{J+1}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-w_{0}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y_{j,t}\right)^{2}+\lambda_{1}\|\widetilde{\bm{w}}\|_{1}+\lambda_{2}\|\widetilde{\bm{w}}\|_{2},$
where $\lambda_{1}>0$ and $\lambda_{2}>0$ are regularization parameters.
4 Identification and Implicit Endogeneity
This section addresses the identification problem of treatment effects and points out the problem of implicit endogeneity in LS-SCMs.
4.1 Linear Models in Expected Outcomes
We revisit the assumptions underlying SCMs. In many existing studies, a linear relationship is assumed between $Y^{N}_{0,t}$ and $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. One possibility is positing the existence of $\bm{w}^{*}\in\Delta^{J}$ (Abadie et al., 2010) such that
$Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}.$
This assumption, often called the perfect pre-treatment fit (Ferman and Pinto, 2021), implies that the stochastic process $\left(Y^{N}_{0,t}\right)_{t\in\mathcal{T}_{0}}$ can be fully replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$ for $t\in\mathcal{T}_{0}$. In linear factor models where $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is often considered unrealistic. Note that this is a specific case and not necessary for $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$.
A more natural assumption is a linear model between $\mathbb{E}_{0,t}[Y^{N}_{0,t}]$ and $(\mathbb{E}_{j,t}[Y^{N}_{j,t}])_{j\in\mathcal{J}^{U}}$.
Assumption 4.1.
There exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$,
$\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}].$ (2)
Compared to Assumption 4.1, the assumption $Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}$ is stronger and less plausible, given that $Y_{j,t}$ is random.
4.2 Asymptotic Bias in Linear Factor Models
We now consider the linear factor model introduced in Definition 2.1, along with Assumption 4.1. This is a specific case of (2). Under the linear factor model, we adopt the following more specialized assumption:
Assumption 4.2.
Under the linear factor model from Definition 2.1, there exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$,
$c_{0}=\sum_{j\in[J]}w^{*}_{j}c_{j}\quad\text{and}\quad\mu_{0}=\sum_{j\in[J]}w^{*}_{j}\mu_{j}\quad\forall t\in\mathcal{T},\ \forall j\in\mathcal{J}.$
Assumption 4.2 is a specific case of Assumption 4.1; if Assumption 4.1 holds, then Assumption 4.2 also holds under the linear factor model in Definition 2.1.
Under these assumptions, Ferman and Pinto (2021) shows that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{0,t}$ suffers from bias and is not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero. We also adopt the following assumption regarding the linear factor model.
Assumption 4.3 (Common and idiosyncratic shocks).
For each $j\in\mathcal{J}$, the linear factor model in Definition 2.1 satisfies the following as $T_{0}\to\infty$:
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\lambda_{t}\xrightarrow{\mathrm{p}}0$.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\xrightarrow{\mathrm{p}}0$.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\lambda^{\top}_{t}\lambda_{t}\xrightarrow{\mathrm{p}}\Omega_{0}$, where $\Omega_{0}$ is positive semidefinite.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\epsilon^{\top}_{j,t}\xrightarrow{\mathrm{p}}\sigma^{2}_{j,\epsilon}I_{J+1}$, where $\sigma^{2}_{j,\epsilon}\geq 0$ is constant.
  • $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\epsilon_{j,t}\lambda_{t}\xrightarrow{\mathrm{p}}0$.
Under Assumptions 4.2 and 4.3, Ferman and Pinto (2021) concludes that the LS-SCM proposed by Abadie and Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$, as the following proposition.
Proposition 4.4 (Proposition 1 in Ferman and Pinto (2021)).
Under Assumptions 4.34.2, we have
$\bm{w}^{\mathrm{Abadie}}\xrightarrow{\mathrm{p}}\overline{\bm{w}}^{\mathrm{Abadie}}\coloneqq\left(\overline{w}^{\mathrm{Abadie}}_{j}\right)_{j\in\mathcal{J}^{U}}$
as $T_{0}\to\infty$, where
$(c_{0},\mu_{0})\neq\left(\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j},\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right),$
unless $\sigma^{2}_{\epsilon}=0$ or $\widetilde{\Phi}\cap\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\{\bm{w}^{\top}\bm{w}\}\neq\emptyset$. Furthermore, for $t\in\mathcal{T}_{1}$,
$\widehat{\tau}^{\mathrm{Abadie}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\lambda_{t}\left(\mu_{0}-\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right)+\left(c_{0}-\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right)+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}\epsilon_{j,t}\overline{w}^{\mathrm{Abadie}}_{j}\right)$
as $T_{0}\to\infty$.
To alleviate this bias problem, Ferman and Pinto (2021) proposes the demeaned SC estimator:
$\widehat{\bm{w}}^{\mathrm{FP}}=\operatorname*{arg\,min}_{\widehat{\bm{w}}^{\mathrm{FP}}\in\Delta^{J}}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left\{\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\right\}^{2},$
where $\overline{Y}^{N}_{j}=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}Y_{j,t}$. They then estimate $\tau_{0,t}$ as
$\widehat{\tau}^{\mathrm{FP}}_{0,t}=Y^{I}_{0,t}-\overline{Y}^{N}_{0}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{FP}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right).$
They show that this approach can reduce the bias seen in the classical LS-SCM estimator. However, there still remains bias in the estimator, which does not vanish even if $T_{0}\to\infty$.
4.3 Implicit Endogeneity
We now interpret the asymptotic bias found by Ferman and Pinto (2021) as an endogeneity problem. We consider (2), that is, $\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}]$, which is more general than the linear factor model.
To illustrate, we consider a simpler case with the least-squares estimator rather than $\bm{w}^{\mathrm{Abadie}}$:
$\widehat{\bm{w}}^{\mathrm{LS}}=\operatorname*{arg\,min}_{\bm{w}\in\mathbb{R}^{J}}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y_{j,t}\right)^{2}.$
We show that $\widehat{\bm{w}}^{\mathrm{LS}}$ is asymptotically biased due to endogeneity. Define $\epsilon_{j,t}=Y_{j,t}-\mathbb{E}_{j,t}[Y_{j,t}]$. Under (2), we obtain
$Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}+\nu_{t},\quad\text{where}\quad\nu_{t}=-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}+\epsilon_{0,t}.$
Noting that $\mathbb{E}[\nu_{t}Y_{j,t}]\neq 0$, we see that there is endogeneity, as the regressor and the error term are correlated. Endogeneity is known to induce bias in least-squares estimation. Let $Q^{*}$ be a $(J\times J)$ matrix whose $(i,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\mathbb{E}_{i,t}[Y_{i,t}]\mathbb{E}_{j,t}[Y_{j,t}]$, and let $\Sigma$ be a $J\times J$ diagonal matrix whose $(j,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\sigma^{2}_{j,t}$.
Theorem 4.5.
As $T_{0}\to\infty$, $\widehat{\bm{w}}^{\mathrm{LS}}\xrightarrow{\mathrm{p}}\overline{\bm{w}}^{\mathrm{LS}}$, where
$\overline{\bm{w}}^{\mathrm{LS}}=\bm{w}^{*}-\left(Q^{*}+\Sigma\right)^{-1}\Sigma\bm{w}^{*}.$
This result follows from Section 5.6 of Greene (2003) and implies that, under (2), there is an asymptotic bias in $\widehat{\bm{w}}^{\mathrm{LS}}$; in other words, least-squares-type estimators do not converge to the true $\bm{w}^{*}$.
Thus, the identification of treatment effects faces a challenge due to implicit endogeneity. While the estimator $\bm{w}^{\mathrm{Abadie}}$ remains standard in many applications, it inherits this endogeneity issue. Such concerns have been noted in prior works (Chernozhukov et al., 2019; Ferman and Pinto, 2021; Ferman, 2021), and Fry (2024) independently frames the bias as an endogeneity problem, proposing an instrumental variable method to address it.
5 SCM by Density Matching
As discussed in Section 4, Assumption 4.1 can be weak or incompatible for identifying $\bm{w}^{*}$ when using least squares. We, therefore, adopt a different assumption for identification. Specifically, we propose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted combination of the densities of $Y^{N}_{j,t}$, and we introduce an SCM based on density matching.
5.1 Identification and Estimation Strategy
Section 4 and the results in Ferman and Pinto (2021) indicate that combining
$\mathbb{E}_{0,t}\left[Y_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]$
with least squares estimators do not yield consistent estimators of $\bm{w}^{*}$. We thus adopt mixture models as a more theoretically favorable but still realistic assumption.
We begin by positing that the distribution of the target unit’s outcome $Y^{N}_{0,t}$ can be expressed as a linear combination of the distributions of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. For simplicity, in this subsection, we assume that $Y^{N}_{j,t}$ is a continuous random variable with a probability density function (PDF). Under this scenario, we suppose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$:
$p^{N}_{0,t}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y)\qquad\forall t\in\mathcal{T}.$ (3)
We next discuss estimating SC weights that satisfy (3). One approach employs optimal transport to minimize the Wasserstein distance (Gunsilius, 2023). In this study, we propose a simpler estimation procedure, as the approach of Gunsilius (2023) involves more complexity than the original SCMs.
Because the mixture model (3) holds, we obtain:
$\mathbb{E}_{0,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{0,t}\right)\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{j,t}\right)\right]\quad\forall t\in\mathcal{T},\ \forall\xi\in\mathbb{R},$
where $\sqrt{-1}$ denotes the imaginary unit. From this, we can estimate $\bm{w}^{*}$ by considering the upper bound of the difference between the expected exponential transforms of $Y^{N}_{0,t}$ and the corresponding weighted sum for $Y^{N}_{j,t}$. Specifically, we have
$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{0,t})\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{j,t})\right]\right|\mathrm{d}\xi$
$=\int_{\xi\in(-h,h)}\Biggl{|}\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right)+\epsilon\left|\xi^{G}\right|\Biggr{|}\mathrm{d}\xi$
$\leq\int_{\xi\in(-h,h)}\Biggl{(}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{|\sqrt{-1}^{\gamma}||\xi|^{\gamma}}{\gamma!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|+\epsilon|\xi^{G}|\Biggr{)}\mathrm{d}\xi$
$=\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}.$
Minimizing
$\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\frac{2\epsilon}{G+1}h^{G+1}$
thus yields an accurate estimate of $\bm{w}^{*}$ that enforces
$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{0,t})\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{j,t})\right]\right|\mathrm{d}\xi=0.$
This approach can be seen as a variant of the method of moments, where we match all moments of $Y^{N}_{0,t}$ with those of the weighted sum of the untreated units.
5.2 Core Assumptions
We now state our core assumption for identifying the SC weights.
Assumption 5.1.
The outcomes $(Y^{N}_{j,t})_{j\in\mathcal{J},t\in\mathcal{T}}$ are bounded.
Assumption 5.2.
There exists $\widetilde{\Phi}^{*}\subset\Delta^{J}$ such that for all $\bm{w}^{*}\in\widetilde{\Phi}^{*}$,
$F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\quad\forall y\in\mathbb{R}.$
Assumption 5.1 ensures that $Y^{N}_{j,t}$ has a finite $G$-th moment for any $G\in\mathbb{N}$. Assumption 5.2 has been employed in earlier work, including Wan et al. (2018), Gunsilius (2023), and Shi et al. (2022). Wan et al. (2018) notes that SCMs effectively rely on Assumption 5.2. Gunsilius (2023) similarly assumes (3) and estimates $\bm{w}^{*}$ by minimizing the Wasserstein distance, while Shi et al. (2022) introduces fine-grained models that, combined with Assumption 5.2, can justify standard SCMs. We provide further discussion of these fine-grained models in Section 8.1.
Although mixture models and linear factor models differ, they are closely related (Shi et al., 2022; Nazaret et al., 2023). In the fine-grained models of Shi et al. (2022), certain implicit assumptions underlie linear factor models, suggesting that mixture model assumptions can also be valid for these settings. Independently, Nazaret et al. (2023) arrive at a similar conclusion. We detail these connections in Section 8.1.
5.3 Estimation of SC Weights
Under the mixture model assumption, we propose estimating SC weights by matching characteristic functions (CHFs). We call our approach the Density Matching SCM (DMSCM).
Let $h\in(0,1)$ be a fixed parameter. We then estimate the SC weights as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\coloneqq\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}\right)_{j\in\mathcal{J}^{U}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|.$ (4)
Using $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$, we can then estimate both the treatment effect and the distributional treatment effect as follows.
5.3.1 Treatment Effect Estimation
Once $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}$ is obtained, we estimate the treatment effect similarly to standard SCMs:
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0},G}Y_{j,t}.$ (5)
We call this estimator the Density Matching SCM (DMSCM) estimator. Algorithm 1 summarizes our proposed estimator and also includes the D2MSCM estimator introduced later.
5.3.2 Distributional Treatment Effect Estimation
To estimate the distributional treatment effect, we can approximate the counterfactual distribution $p_{0,t}(y)$ via bootstrap. Specifically, we resample $L>1$ observations $\{Y^{l,\dagger}_{0}\}_{l=1}^{L}$ from $\{Y_{0,t},Y_{1,t},\dots,Y_{J,t}\}_{t\in\mathcal{T}_{1}}$ as follows: for each $l\in\mathcal{L}\coloneqq\{1,2,\dots,L\}$,
  1. 1.
    For each $j\in\mathcal{J}$, resample $Y^{l,\dagger}_{j}$ from $\{Y_{j,t}\}_{t\in\mathcal{T}_{1}}$.
  2. 2.
    Choose $Y^{l,\dagger}_{0}$ from $\left(Y^{l,\dagger}_{1},\dots,Y^{l,\dagger}_{J}\right)$ with probability $\left(\widehat{w}^{\mathrm{DMSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$.
Because our primary focus is often on $\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p_{0,t}(y)$, this bootstrap procedure enables us to approximate the distribution of the counterfactual outcomes.
Algorithm 1 DMSCM.
  Parameter: The number of moments $G$.
  Obtain trained weights $\widehat{\bm{w}}^{\mathrm{DMSCM}}$ as (4).
  Estimate the treatment effect as (5).
6 Convergence Analysis
We now provide consistency results for our proposed SC weight estimator. The following theorem presents these results, including consistency for the treatment effect $\tau_{0}$.
Assumption 6.1.
Assume that the density function of $\varepsilon_{j,t}$ exists and is bounded for each $j\in\mathcal{J}$ and $t\in\mathcal{T}_{0}$, and that the stochastic process $\{\varepsilon_{j,t}\}_{t=1}^{T_{0}}$ is stationary, strongly mixing, with the sum of mixing coefficients bounded by $M$.
Theorem 6.2.
Suppose that Assumptions 5.15.2 and 6.1 hold. Fix $h\in(0,1)$. Define
$\widetilde{\Phi}^{\dagger}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|.$
Then, it holds that
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}\quad\text{as }T_{0}\to\infty.$
Additionally, $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$, and
$\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}\xrightarrow{\mathrm{p}}\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)\quad\text{as }T_{0}\to\infty\text{ and }G\to\infty.$
The proof is shown in Appendix A. Note that this result also applies to $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ as a special case. Consequently, under the respective assumptions, $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ and $\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}$ are asymptotically unbiased.
7 Inference
This section presents methods for statistical inference using our proposed DMSCM estimator.
7.1 Conformal Inference on the Treatment Effect
We consider statistical inference for the treatment effect using our proposed DMSCM and D2SCM, focusing on testing a sharp null hypothesis, $H_{0}:\tau_{0,t}=\alpha_{t}$ for some $t\in\mathcal{T}_{t}$ and $\bm{\alpha}=(\alpha_{t})_{t\in\mathcal{T}_{1}}$. Specifically, we adopt the conformal inference approach of Chernozhukov et al. (2021), which has been used by, for example, Ben-Michael et al. (2021) and Ferman and Pinto (2021).
Our goal is to construct a confidence interval whose $p$-value $\widehat{p}$ satisfies
$\mathrm{Pr}\left(\widehat{p}\leq a\right)\to a\quad\text{for all }a\in(0,1)\quad\text{as }T_{0}\to\infty\text{ and }T_{1}\text{ is fixed}.$ (6)
To achieve this, conformal inference proceeds as follows. First, for the sharp null hypothesis
$H_{0}:\tau_{0,t}=\alpha_{t},$
we create adjusted post-treatment outcomes for the treated unit $Y^{N\sharp}_{0,t}=Y^{I}_{0,t}-\alpha_{t}$ and combine them with the original dataset as
$\left(Y_{0,1},\dots,Y_{0,T_{0}},Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right).$
We then apply our proposed method to this augmented dataset (comprising
$\left(Y_{0,1},\dots,Y^{N\sharp}_{0,(T_{0}+1)},\dots,Y^{N\sharp}_{0,T}\right)$
and $(Y_{j,1})_{t\in\mathcal{T},j\in\mathcal{J}^{U}}$) to obtain adjusted weights $\widehat{\bm{w}}(\alpha)$ and an adjusted predictor of the counterfactual outcome, denoted $\widehat{Y}^{N}_{0,t}(\bm{\alpha})$.
Finally, we compute the $p$-value by assessing whether the adjusted residual “conforms” with the pre-treatment residuals. Define $u_{t}(\bm{\alpha})=Y_{0,t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{0}$, and $u_{t}(\bm{\alpha})=Y^{I}_{0,t}-\alpha_{t}-\widehat{Y}^{N}_{0,t}(\bm{\alpha})$ for $t\in\mathcal{T}_{1}$. For $j\in\mathcal{T}$, let $\pi_{j}$ be a permutation of indices defined by $\pi_{j}(t)=t+j$ if $t+j\leq T$, and $\pi_{j}(t)=t+j-T$ otherwise. Let $\Pi=\{\pi_{1},\dots,\pi_{T-1}\}$. Then the $p$-value is defined as $\widehat{p}(\bm{\alpha})=1-\widehat{F}(\bm{\alpha})$, where
$\widehat{F}(\bm{\alpha})=\frac{1}{T}\sum_{j\in\mathcal{T}}\mathbbm{1}\left[\frac{1}{T}\sum_{t\in\mathcal{T}}\left|u_{t}(\bm{\alpha})\right|<\frac{1}{T}\sum_{\pi_{j}\in\mathcal{T}}\left|u_{\pi_{j}}(\bm{\alpha})\right|\right].$
Because $Y^{N}_{0,t}$ is random, inverting this test yields a confidence interval for $\tau_{0,t}$, which is equivalent to constructing a conformal prediction set (Vovk et al., 2005) for $Y^{N}_{0,T}$. The $(1-\xi)$ confidence interval is
$\widehat{C}^{\mathrm{conf}}_{1-\xi}=\left\{\bm{\alpha}\in\mathcal{A}\bigm{|}\widehat{p}(\bm{\alpha})>\xi\right\},$
where $\mathcal{A}$ is a set of candidate values. Theorem 1 of Chernozhukov et al. (2021) guarantees (6).
7.2 Two-Sample Homogeneity Test
To test distributional treatment effects, consider the null and alternative hypotheses $H_{0}:p^{I}_{t}=p^{N}_{t}$ versus $H_{1}:p^{I}_{t}\neq p^{N}_{t}$. One can perform a two-sample homogeneity test in combination with our distributional treatment effect estimator. For instance, one may employ maximum mean discrepancy (MMD) (Gretton et al., 2012) to implement this hypothesis test.
8 Discussion
We conclude by discussing some remaining issues.
8.1 Justification of Mixture Models
To clarify assumptions in SCMs, Shi et al. (2022) proposes fine-grained models wherein each unit is composed of multiple components, and the outcomes of treated and untreated units are averages of the outcomes across these components. For example, in the empirical study of California’s tobacco control program, the outcome of interest (annual per capita cigarette consumption) at the state level is an average of individuals’ consumption in the state. Building on this framework, one can justify mixture model assumptions by relating them to linear factor models.
Assume there are $U_{j}$ components in unit $j\in\mathcal{J}$. For each $u\in\mathcal{U}_{j}\coloneqq\{1,2,\dots,U_{j}\}$, let $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ be the potential outcomes of component $u$ in unit $j$ at time $t$. Suppose there is a $d_{W}$-dimensional unobserved random variable $W_{j,t,u}\in\mathbb{Z}^{d_{W}}$ capturing individual characteristics (e.g., age or income). We assume $(Y^{I}_{j,t,u},Y^{N}_{j,t,u},W_{j,t,u})$ is i.i.d. across $u\in\mathcal{U}_{j}$.
The relationship between $W_{j,t,u}$ and potential outcomes $(Y^{I}_{j,t,u},Y^{N}_{j,t,u})$ can be linear or nonlinear, and possibly time-varying. Let $p_{j,t}(w,y)$ be the PDF of $(W_{j,t},Y^{N}_{j,t})$, and let $p_{j,t}(y\mid w)$ be the conditional PDF of $Y^{N}_{j,t}$ given $W_{j,t}$, while $p_{j,t}$ is the marginal PDF of $W_{j,t}$. We assume that the intervention is administered at a group level and that individuals within each group comply with their group-level treatment.
Under these conditions, Shi et al. (2022) shows that the following assumptions are required for the linearity of SCMs in (2):
Assumption 8.1 (Independent Causal Mechanism; ICM).
Conditional on the causes $W_{j,t,u}$, the potential outcome $Y^{N}_{j,t,u}$ is independent of the unit $j\in\mathcal{J}$. For the population distribution $j\in\mathcal{J}$ at time $t\leq\mathcal{T}$, the joint distribution of $(Y^{N}_{j,t,u},W_{j,t,u})$ is given by $p_{j,t}(w,y)=r_{j,t}(w)p_{t}(y\mid w)$.
Assumption 8.2 (Stable distributions).
Decompose the causes as $W_{j,t,u}=\{U_{j,t,u},S_{j,t,u}\}$. Let $S$ denote the subset of causes that differentiates the treated unit from the selected untreated units (the minimal invariant set). We assume that, for all groups, the distribution of $S$ does not change for all time periods $t\leq T$. Hence, $r_{j,t}(w)=\zeta_{j}(s)\xi_{t}(u\mid s)$, where $\zeta_{j}(s)$ is the PDF of $S_{j,t,u}$, and $\xi_{t}(u\mid s)$ is the PDF of $U_{j,t,u}$ given $S_{j,t,u}$.
Assumption 8.3 (Sufficiently similar untreated units).
Let $S$ be the minimal invariant set for the target and untreated units. The untreated units are sufficiently similar if the cardinality of this minimal invariant set satisfies $\left|\mathcal{J}^{U}\right|\geq\lvert S\rvert$.
Assumption 8.4 (Common support).
Let $(s_{1},\dots,s_{R})$ be the support of $S$. For each $s$, there is at least one untreated unit $j$ with $\zeta_{j}(S=s)>0$.
Proposition 8.5 (Causal identifiability. From Theorem 1 of Shi et al. (2022)).
Suppose Assumptions 8.18.4 hold. Then there exist weights $\{w^{*}_{j}\}_{j\in[J]}$ such that $p^{N}_{0,t}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y)$ holds.
This result also implies that the mixture model (3) holds (Nazaret et al., 2023), aligning it with linear factor models. Hence, we obtain the following corollary for linear factor models similar to Ferman and Pinto (2021):
Corollary 8.6.
Suppose that Assumptions 5.1, 6.1, and 8.18.4 hold. Then
$\bm{w}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\quad\text{as }T_{0}\to\infty,$
and $\widehat{\tau}^{\mathrm{DMSCM}}_{0,t}$ converges to
$\tau_{0,t}+\left(\epsilon_{0,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}\right)$
in probability as $T_{0}\to\infty$ and $G\to\infty$.
Proof Sketch. From Proposition 8.5, under Assumptions 8.18.4, (3) holds. Therefore, applying Theorem 6.2 yields the desired result under (1) with $c_{j}=0$ for each $j\in\mathcal{J}$.
8.2 Auxiliary Covariates
We can also extend our method to incorporate auxiliary covariates if they are observable. For each unit $j\in\mathcal{J}$, let
$X_{j,t}\coloneqq(X_{j,t,1},X_{j,t,2},\dots,X_{j,t,K})\in\mathbb{R}^{K}$
be $K$-dimensional covariates that are unaffected by treatment. Define
^$\widehat{\ell}_{k}(\bm{w})=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(X_{0,t,k}-\sum_{j\in\mathcal{J}^{U}}w_{j}X_{j,t,k}\right).$
When these covariates are available, we can extend the DMSCM estimator as
$\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0}}\coloneqq v_{0}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\sum_{k=1}^{K}v_{k}\widehat{\ell}^{2}_{k}(\bm{w}),$
where $\bm{v}\coloneqq(v_{k})_{k=0}^{K}$ are weights such that $v_{k}\geq 0$ and $\sum_{k=0}^{K}v_{k}=1$.
8.3 Distributional SCMs and Optimal Transport
Gunsilius (2023) propose distributional SCMs, which assume
$F^{-1}_{Y_{0,t},N}(q)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F^{-1}_{Y_{j,t},N}(q)\quad\forall q\in(0,1),$
where $F^{-1}_{Y_{j,t},N}$ is the quantile function defined by
$F^{-1}_{Y_{j,t},N}(q)\coloneqq\inf\left\{y\in\mathbb{R}\colon F_{Y_{j,t},N}(y)\geq q\right\}.$
To estimate $\bm{w}^{*}$, Gunsilius (2023) minimize the Wasserstein distance, referring to their method as DiSCo. Concretely, the DiSCo estimator is
$\widehat{\bm{w}}^{\mathrm{DiSCo}}\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\frac{1}{M}\sum_{m=1}^{M}\left|\widehat{F}^{-1}_{Y^{N}_{0,t}}(V_{m})-\sum_{j\in\mathcal{J}^{U}}w_{j}\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})\right|,$
where $\widehat{F}^{-1}_{Y^{N}_{j,t}}(V_{m})$ is the empirical quantile function of $Y^{N}_{j,t}$, and $\{V_{m}\}_{m=1}^{M}$ are i.i.d. draws from the uniform distribution on $[0,1]$.
8.4 Demeaned Density Matching SCM (D2MSCM)
Finally, we consider the case where the target distribution does not precisely satisfy (3), yet any deviation disappears after demeaning. Concretely, instead of (3), we assume
$p^{N}_{0,t}(y-c_{0})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p^{N}_{j,t}(y-c_{j})\qquad\forall t\in\mathcal{T}.$
From the perspective of the linear factor model (1) and fine-grained models (Shi et al., 2022), this assumption permits $c_{j}\neq 0$ for some $j\in\mathcal{J}$. Under the previous mixture model, we implicitly required $c_{j}=0$ for all $j\in\mathcal{J}$, which may be restrictive in practice.
To estimate the SC weights in this setting, we first estimate $c_{j}$ by taking the average of $(Y_{j,t})_{t\in\mathcal{T}_{0}}$ and subtracting it from each $Y_{j,t}$. This demeaned approach is inspired by the proposal of Ferman and Pinto (2021) but is applied here to density matching. Let $h\in(0,1)$ be a fixed parameter. We estimate the weights as
$\widehat{\bm{w}}^{\mathrm{D2MSCM}}_{T_{0}}$ $\coloneqq\left(\widehat{w}^{\mathrm{D2MSCM}}_{j,T_{0}}\right)_{j\in\mathcal{J}^{U}}$
$\coloneqq\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{\gamma=1}^{G}\frac{2h^{\gamma+1}}{(\gamma+1)!}\biggl{|}\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}-\overline{Y}^{N}_{0}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)^{\gamma}\right)\biggr{|}.$
We call this procedure demeaned DMSCM (D2SCM). Using these weights, we can estimate the treatment effect as
$\widehat{\tau}^{\mathrm{D2MSCM}}_{0,t}\coloneqq Y^{I}_{0,t}-\biggl{(}\overline{Y}^{N}_{0}+\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{D2MSCM}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)\biggr{)}.$
Figure 2: Results of simulation studies.
Figure 3: Results of simulation studies on ATE prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the mean squared prediction error.
Figure 4: Results of simulation studies on distributional treatment effect prediction. The $x$-axis is the number of untreated units $J$, and the $y$-axis is the MMD.
9 Experiments
We conduct experiments using both simulated and real-world data. Additional results can be found in Appendix 9.1.
9.1 Simulation Studies
We first perform simulation studies to compare our proposed SCMs with traditional SCMs. Set $J\in\{10,30,60\}$ and $G\in\{2,3,5,10,20,30,40,50,60,70,80,90,100\}$, $T_{0}=30$, $T_{1}=100$, and $K=5$.
We consider SCMs for
$p_{0,t}(\bm{z})=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}p_{j,t}(\bm{z}),$
where $\bm{z}$ is a $(K+1)$-dimensional (i.e., $6$-dimensional) vector: the first element corresponds to $Y_{j,t}$ and the remaining elements correspond to $X_{j,t}$. For each $j\in\mathcal{J}$, let $p_{j,t}$ be the density of a multivariate normal distribution
$\mathcal{N}\left(\bm{\mu}_{j,t},\bm{\sigma}^{2}_{j,t}\right),$
where $\bm{\mu}_{j,t}=(\mu_{j,t,1},\dots,\mu_{j,t,6})$ and $\bm{\sigma}^{2}_{j,t}$ is a $(6\times 6)$ diagonal matrix whose $(k,k)$-element is $\sigma^{2}_{j,t,k}$. Each $\mu_{j,t,k}$ is drawn from a standard normal distribution, and each $\sigma^{2}_{j,t,k}$ is drawn from a uniform distribution on $[1,20]$. For each $j\in\mathcal{J}$, we modify $\mu_{j,t,1}$ and $\sigma^{2}_{j,t,1}$ by adding noise:
$\mu_{j,t,1}\leftarrow\mu_{j,t,1}+\epsilon_{j,t},\quad\sigma^{2}_{j,t,1}\leftarrow\sigma^{2}_{j,t,1}+\widetilde{\epsilon}_{j,t},$
where $\epsilon_{j,t}$ and $\widetilde{\epsilon}_{j,t}$ follow $\mathcal{N}(0,10)$. If $\widetilde{\epsilon}_{j,t}$ is less than or equal to $0.1$, we set $\widetilde{\epsilon}_{j,t}=0.1$ to avoid near-zero variance.
For the treated unit, we let $Y^{N}_{0,t}\sim p_{0,t}(\bm{z})$ for $t\in[t_{0},T_{1}]$ and set $Y^{I}_{0,t}=20+Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$, implying $\tau_{0,t}=20$. We draw $\bm{w}^{*}$ from a uniform distribution over $[0,1]^{J}$ and then normalize it so that its entries sum to $1$.
We apply DMSCM and the method of Abadie and Gardeazabal (2003) (referred to here as Abadie) $1000$ times and evaluate the estimation error of the treatment effect. We plot the estimation error box plots in Figure 3, indicate that our proposed method achieves smaller estimation errors. Moreover, the estimation error declines as $G$ increases.
Next, we consider another set of simulation studies under conditions similar to those above, but with the following changes: $J\in\{1,5,10,15,20,\dots,35,40,45,50\}$, $G\in\{2,3,5,10\}$, $T_{0}=30$, $T_{1}=1000$, and $K=5$. All other settings remain as in Section 9.1.
We again apply DMSCM and Abadie $1000$ times and compute the error of the treatment effect estimation (Figure 3). We focus on how the mean squared estimation error changes as $J$ grows. While the estimation error under the method of Abadie (2002) grows with $J$, our proposed method does not exhibit such growth. Additionally, our method maintains lower estimation errors than the method of Abadie (2002).
A similar pattern appears in distributional treatment effect estimation. We measure the discrepancy between the true and estimated distributions using the Maximum Mean Discrepancy (MMD, Gretton et al., 2012), a probability metric based on kernel embeddings. In the top panel of Figure 4, we show how MMD varies with $J$. As before, our proposed methods outperform the method of Abadie (2002) in this setting. In the bottom panel of Figure 4, we plot MMD for $G\in\{2,10,50,100\}$. As $G$ grows (i.e., as more moment conditions are imposed), the predictive performance improves.
Our observations may connect with findings by Spiess et al. (2023), who examine high-dimensional SCMs. Exploring these connections will be part of our future research.
Figure 5: Counterfactual outcomes in empirical analysis.
Figure 6: Counterfactual distributions in empirical analysis.
Table 2: Results of treatment effect estimation and conformal inference. Estimated treatment effect $\widehat{\tau}_{0,t}$ and $90\%$ confidence interval $(\underline{\tau}_{0,t},\overline{\tau}_{0,t})$.
Terrorist conflict in Basque country.
Abadie DMSCM
$\gamma$- $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$ $-0.800$ $-0.845$ $-1.054$ $-0.279$ $-0.138$
$\overline{\tau}_{0,t}$ $0.522$ $0.699$ $0.495$ $0.521$ $0.578$
$\underline{\tau}_{0,t}$ $-2.587$ $-2.150$ $-2.605$ $-2.616$ $-2.590$
Tobacco control program in California.
Abadie DMSCM
$\gamma$- $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$ $-26.148$ $-27.744$ $-32.306$ $-35.080$ $-36.879$
$\overline{\tau}_{0,t}$ $8.698$ $7.852$ $8.898$ $9.544$ $11.748$
$\underline{\tau}_{0,t}$ $-36.798$ $-29.268$ $-34.783$ $-37.308$ $-35.243$
Reunification of Germany.
Abadie DMSCM
$\gamma$- $2$ $5$ $10$ $100$
$\widehat{\tau}_{0,t}$ $-1492.374$ $-1253.616$ $-1419.120$ $-1324.448$ $-510.765$
$\overline{\tau}_{0,t}$ $2635.583$ $-1008.822$ $-591.420$ $-716.376$ $2327.944$
$\underline{\tau}_{0,t}$ $-2390.413$ $-2738.232$ $-1989.323$ $-2018.878$ $-2625.129$
9.2 Empirical Analysis
We now present empirical studies based on Abadie and Gardeazabal (2003), Abadie et al. (2010), and Abadie et al. (2015).
Terrorist conflict in the Basque Country. Abadie and Gardeazabal (2003) investigate the effect of terrorist conflict in the Basque Country on gross domestic product (GDP). In this example, $J=16$, $K=14$, and $(t_{0},T_{0},T_{1})=(1955,1970,1997)$. Basque Euskadi Ta Askatasuna (Basque Homeland and Liberty; ETA) was founded in 1959 as a Basque separatist organization in Spain, seeking an independent Basque state primarily through terrorist activities in the Basque Country region. Abadie and Gardeazabal (2003) use annual Spanish regional-level panel data from 1955 to 1997, with a particular focus on the 1970s when ETA’s terrorist activity was significant. The dataset includes fifteen years of pre-intervention observations. The set of untreated units consists of $17$ other Spanish regions. Each region’s record contains the outcome of interest (GDP per capita) and additional covariates (investment rate, population density, five sectoral output measures, and four measures of human capital), all averaged from 1955 to 1968.
Tobacco control program in California. Abadie et al. (2010) study the effect of a large tobacco control program adopted in California in 1988 (described in the Introduction). Here, $J=20$, $K=5$, and $(t_{0},T_{0},T_{1})=(1970,1989,2000)$.
Reunification of Germany. Abadie et al. (2015) examine the impact of Germany’s reunification in 1990. Because no single country provided a suitable comparison, SCMs were used to create a composite unit. In this example, $J=8$, $K=31$, and $(t_{0},T_{0},T_{1})=(1960,1990,2003)$.
Separately, Abadie et al. (2010) also investigate the causal effect of Proposition 99, a comprehensive tobacco control program adopted by California. They use annual state-level panel data from 1970 to 2000. The program was implemented in January 1989, providing eighteen years of pre-intervention observations. Four states that adopted large-scale tobacco control programs (Massachusetts, Arizona, Oregon, and Florida) and states that raised cigarette taxes by at least \$0.50 between 1989 and 2000 (Alaska, Hawaii, Maryland, Michigan, New Jersey, New York, and Washington) as well as the District of Columbia are excluded from the set of 38 control states. The outcome of interest is annual per capita cigarette sales in packs; covariates include the average retail price of cigarettes, log of state income per capita, the percentage of the population aged 15–24, and per capita beer consumption, all averaged between 1980 and 1988.
Results. Table 2 reports the estimated treatment effects and 90% confidence intervals. Figures 6 and 6 show the estimated treatment effects and the distributional treatment effects, respectively.
In each dataset, we plot the counterfactual outcomes and distributions in Figures 6 and 6. We also perform conformal inference on the sharp null hypothesis $H_{0}:\tau_{0,(T_{0}+1)}=\cdots=\tau_{0,T_{1}}=0$ for the treatment effect. For both Abadie’s method and DMSCM, we present the 90% confidence interval of the treatment effect in Table 2. Our proposed method fits the observed outcomes well in the pre-treatment period ($t\in\mathcal{T}_{0}$) and, accordingly, produces plausible counterfactual outcomes and distributions. In the conformal inference analysis, our method does not reject the null hypothesis in the same manner as Abadie’s estimator, except for the DMSCM in the California tobacco control example.
10 Conclusion
We have highlighted the problem of implicit endogeneity in the linear regression models commonly employed by existing SCMs. To address this issue, we proposed assuming mixture models among outcome distributions and developed a density-matching SCM. By matching densities, our method consistently estimates SC weights and reinforces the theoretical foundation for using SCMs in comparative studies. We demonstrated the effectiveness of this approach through both simulation experiments and empirical applications.
Appendix A Proof of Theorem 6.2
First, we confirm the basic facts about the uniqueness of the CHF.
Proposition A.1 (Theorem 26.2 in Billingsley (1995)).
Distinct measures cannot have the same characteristic function.
Let us define
$\widehat{Q}_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right)\right|,$
$Q_{G}(\bm{w})\coloneqq\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|.$
Proof.
First, from the definition of $\Delta^{J}$, $\Delta^{J}$ is a compact parameter space. Next, $\widehat{Q}(\bm{w})$ converges uniformly to $Q(\bm{w})$ as $T_{0}\to\infty$ under Assumption 6.1. Furthermore, $Q(\bm{w})$ is continuous with respect to $\bm{w}$.
Under these conditions, for any $\varepsilon>0$, we have with probability approaching one, we have
  1. 1.
    $\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}(\bm{w}_{0})-\varepsilon/3$;
  2. 2.
    $\widehat{Q}_{G}(\widehat{\bm{w}}^{\mathrm{DMSCM}})>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3$;
  3. 3.
    $\widehat{Q}_{G}\left(\bm{w}_{0}\right)>Q_{G}(\bm{w}_{0})-\varepsilon/3$.
Therefore, with probability approaching one,
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\widehat{Q}_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)-\varepsilon/3>\widehat{Q}_{G}\left(\bm{w}_{0}\right)-2\varepsilon/3>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon.$
Thus, for any $\varepsilon>0$, $Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>Q_{G}\left(\bm{w}_{0}\right)-\varepsilon$ with probability approaching one. Let $\mathcal{N}$ be any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Since $\Theta\cap\mathcal{N}^{c}$ is compact, and $Q_{G}(\bm{w})$ is continuous with respect to $\bm{w}$, we have
$\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})=Q_{G}(\bm{w}^{\dagger})<Q_{G}(\bm{w}_{0})$
for some $\bm{w}^{\dagger}\in\Theta\cap\mathcal{N}^{c}$. Thus, choosing $\varepsilon=Q_{G}(\bm{w}_{0})-\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w})$, it follows that
$Q_{G}\left(\widehat{\bm{w}}^{\mathrm{DMSCM}}\right)>\sup_{\bm{w}\in\Delta^{J}\cap\mathcal{N}^{c}}Q_{G}(\bm{w}),$
and hence $\widehat{\bm{w}}^{\mathrm{DMSCM}}\in\mathcal{N}$, where recall that $\mathcal{N}$ is any open subset of $\Delta^{J}$ containing $\widetilde{\Phi}^{*}$. Thus, $\widehat{\bm{w}}^{\mathrm{DMSCM}}_{T_{0},G}\xrightarrow{\mathrm{p}}\bm{w}^{*}_{G}\in\widetilde{\Phi}^{\dagger}$ holds as $T_{0}\to\infty$. Additionally, from Proposition A.1, we have $\bm{w}^{*}_{G}\to\bm{w}^{*}\in\widetilde{\Phi}^{*}$ as $G\to\infty$. ∎